### v1

# DeFiScope: Detecting Various DeFi Price Manipulations with LLM Reasoning

Abstract.
DeFi (Decentralized Finance) is one of the most important applications of today’s cryptocurrencies and smart contracts. It manages hundreds of billions in Total Value Locked (TVL) on-chain, yet it remains susceptible to common DeFi price manipulation attacks. Despite state-of-the-art (SOTA) systems like DeFiRanger and DeFort, we found that they are less effective to non-standard price models in custom DeFi protocols, which account for 44.2% of the 95 DeFi price manipulation attacks reported over the past three years.
In this paper, we introduce the first LLM-based approach, DeFiScope, for detecting DeFi price manipulation attacks in both standard and custom price models. Our insight is that large language models (LLMs) have certain intelligence to abstract price calculation from code and infer the trend of token price changes based on the extracted price models. To further strengthen LLMs in this aspect, we leverage Foundry to synthesize on-chain data and use it to fine-tune a DeFi price-specific LLM. Together with the high-level DeFi operations recovered from low-level transaction data, DeFiScope detects various DeFi price manipulations according to systematically mined patterns. Experimental results show that DeFiScope achieves a high precision of 96% and a recall rate of 80%, significantly outperforming SOTA approaches. Moreover, we evaluate DeFiScope’s cost-effectiveness and demonstrate its practicality by helping our industry partner confirm 147 real-world price manipulation attacks, including discovering 81 previously unknown historical incidents.
copyright: none
1. Introduction
DeFi represents a form of finance that eliminates traditional financial intermediaries by utilizing smart contracts on a blockchain. A smart contract (Zou et al., 2019) is a self-executing program with the terms of the agreement between the buyer and the seller directly written into lines of code. The code and the agreements contained therein exist across a distributed, decentralized blockchain network (Dinh et al., 2018). Smart contracts enable trusted transactions and agreements to be carried out among disparate, anonymous parties without the need for a central authority, legal system, or external enforcement mechanism.
While price manipulation is a pervasive risk in traditional financial markets, it poses even more significant damage in the context of DeFi. In the cryptocurrency space, attackers strategically add or remove large amounts of tokens in specific contracts, or impact the total supply of target tokens by burning or minting, to exploit vulnerabilities in the pricing mechanisms of DeFi applications, ultimately inflating or deflating the price of cryptocurrencies (Cointelegraph, 2024). To detect price manipulation attacks in DeFi applications, researchers have proposed several state-of-the-art (SOTA) approaches, primarily focusing on transaction monitoring-based methods that offer real-time protection, e.g., DeFiRanger (Wu et al., 2023) and DeFort (Xie et al., 2024).
However, our analysis indicates that these SOTA approaches are less effective to non-standard price models in custom DeFi protocols. This is because they typically require token exchange rates to identify abnormal price changes, which are only suitable to calculate under standard price models such as CPMM (Constant Product Market Makers) and Stableswap Invariant (detailed in §2.2). Unfortunately, our subsequent evaluation shows that 44.2% of the 95 DeFi price manipulation attacks reported in the past three years used non-standard price models. Hence, instead of explicitly calculating the exchange rate for a pair of tokens, we aim to capture only the abnormal price fluctuations of tokens, which can be directly derived from the high-level price model and changes in token balances.
To achieve this objective, we sought help from large language models (LLMs), considering that their trained intelligence might aid in inferring price changes associated with DeFi operations during the transaction process. As such, we introduce DeFiScope, the first LLM-based approach to detecting DeFi price manipulation attacks in both standard and custom price models. DeFiScope features several novel designs, such as constructing a transfer graph to recover high-level DeFi operations (§5) and systematically mining price manipulation patterns across all types of DeFi applications (§6). Among them, the key design is to fine-tune a DeFi price-specific LLM (§4), where we propose (i) simulating transactions using Foundry (Github, 2024) to generate our own fine-tuning data, and (ii) conducting a Chain-of-Thought (CoT)-style fine-tuning that integrates both on-chain data and the price context. The ablation study shows that fine-tuning increases the overall detection effectiveness by up to 31% compared to the baseline LLM (under DeFiScope) without fine-tuning.
To thoroughly evaluate DeFiScope’s effectiveness and practicality, we conduct both a benchmark and a large-scale experiment. In the benchmark experiment, we collected 95 real-world price manipulation attacks from multiple sources and compared DeFiScope with three SOTA tools. The results show that DeFiScope achieves a superior detection accuracy (recall) of 80%, significantly higher than 51.6% with DeFiRanger, 52.6% with DeFort, and 35.8% with DeFiTainter (Kong et al., 2023) (a static source code analysis tool). In a large-scale experiment involving 968 suspicious transactions (a mix of various DeFi attacks, not just price manipulations) from our industry partner and 96,800 benign transactions from DeFort (Xie et al., 2024), we helped confirm 147 out of 153 price manipulation attacks reported by DeFiScope, yielding a precision of 96%. In particular, 81 of these attacks are previously unknown historical incidents. Moreover, DeFiScope achieves zero false alarms on benign transactions, with an average of 2.5-second time overhead across suspicious and benign transactions and merely \$0.0107 per-request LLM inference cost.
In sum, this paper makes the following contributions:
  • We introduce the first LLM-based approach, DeFiScope, for automatic on-chain price manipulation attack detection. In particular, LLMs, when properly fine-tuned, have the capability to abstract price models and infer price changes.
  • To support DeFiScope’s detection, we propose a graph-based method to recover high-level DeFi operations and systematically mine eight price manipulation patterns.
  • We extensively evaluate DeFiScope with three real-world transaction datasets, showing DeFiScope’s superior performance over SOTA tools in terms of precision and recall.
Open Science and Ethics. To facilitate future research, all source code and experimental results are available in an (anonymous) GitHub repository (Anonymous, 2024). For ethical research, we ensure that all attacks elaborated in this work have been published or investigated by authoritative security companies through their established channels, and did not involve any undisclosed zero-day vulnerabilities.
2. Background and Motivation
As Figure 1 shows, on blockchains, two primary account types exist, i.e., externally owned accounts (EOAs), controlled by individuals using private keys, and contract accounts (CAs), governed by their contract code. External transactions are initiated externally by EOAs, while internal transactions occurs when a smart contract calls another contract internally. Smart contracts exchange information with other smart contracts through internal transactions, where a transaction involving communication between contracts can be regarded as a sequence of function calls, which has been greatly utilized in DeFi protocols for interoperability. For simplicity, we denote as user invocation an internal transaction between the user-controlled smart contracts and other smart contracts, which plays a key role in price manipulation attacks.
2.1. DeFi Protocols
A DeFi protocol typically consists of multiple smart contracts and offers various services including trading on decentralized exchange, lending, and yield farming.
Decentralized EXchange (DEX). A DEX is a peer-to-peer marketplace where users can trade cryptocurrencies. The two predominant forms of DEXs are order book DEXs and automated market makers (AMMs). In the order book DEXs, the management and matching processes of open buy and sell orders take place off-chain while the settlement of trades occurs on-chain. Alternatively, the AMM is fully on-chain, and maintains liquidity through liquidity pools without replying on real-time buyers and sellers. A liquidity pool maintains two or more tokens, and the prices of these tokens are determined by a constant mathematical formula. In this paper, we only target fully on-chain DEXs using AMM (Chainlink, 2024; coinbase, 2024).
Lending. To borrow assets from a lending protocol, borrowers are required to deposit other cryptocurrencies as collateral first. To ensure that borrowers can repay their loans, the collateral factor of supplied assets is utilized to calculate the maximum value of assets that users can borrow. For instance, consider a scenario where the collateral factor of ETH is set to 0.75, borrowers can only obtain assets up to 75% of the value of ETH they have deposited as collateral. If the value of the collateral decreases and falls below a certain level (the health factor of a borrower is less than 1), the liquidation of the collateral (the lending protocol sells the collateral at a discount rate) occurs to repay the borrowed amount (moai finance, 2024).
Yield Farming. It provides a way to earn rewards in the form of Annual Percentage Yields (APYs). A yield farmer could deposit cryptocurrencies into yield-farming applications, which invest these assets according to various strategies. Based on the returns from these investments, yield farmers can obtain rewards. There are two types of yield farming. The first involves users depositing cryptocurrencies into an application to obtain proof tokens, which can be traded, allowing holders to withdraw the corresponding deposited assets. The second, a stake-and-claim model, has users lock their assets into the contract and later claim them back. This process does not generate proof tokens, meaning the staked shares are non-tradable, and only the staking account can make the claim.
Figure 1. The blockchain transactions.
2.2. Price Models
A price model represents the pricing mechanism within the DeFi application, which are typically expressed as equations correlating the price of a certain token with the balances of and the total supply of various tokens as well as other constant.
Constant Product Market Maker (CPMM). CPMM is one of the most prevalent DeFi AMMs and used in many well-known DEXs (Uniswap, 2024; PancakeSwap, 2024). It maintains liquidity through a constant function $R_{x}*R_{y}=k$, where $k$ is a constant, and $R_{x}$ and $R_{y}$ represent the reserves of two distinct assets $token_{x}$ and $token_{y}$ in a liquidity pool. In a swap operation, let $\Delta x$ amount of $token_{x}$ can exchange for $\Delta y$ amount of $token_{y}$, resulting in $(x+\Delta x)*(y-\Delta y)=k$ where the instantaneous price of $token_{x}$ denominated in $token_{y}$ is $P_{x,y}=\frac{y}{x}$ (Wu et al., 2023; Gogol et al., 2024).
Stableswap Invariant. The Stableswap Invariant (Curve.fi, 2024c), widely used in Curve AMMs (Curve.fi, 2024a), and is defined as follows:
$\frac{a\prod_{i=1}^{n}x_{i}}{(D/n)^{n}}\cdot D^{n-1}\cdot\sum_{i=1}^{n}x_{i}+\prod_{i=1}^{n}x_{i}$ $=\frac{a\prod_{i=1}^{n}x_{i}}{(D/n)^{n}}\cdot D^{n}+\frac{D^{n}}{n^{n}}$
where $a$ indicates a constant amplification coefficient and n is the number of tokens in the liquidity pool, while $D$ represents the total underlying value, i.e., the total amount of tokens in the pool when token prices are equal, and $x_{i}$ denotes the current reserve of $token_{i}$. As $a$ approaches zero, the function behaves similarly to a CPMM.
Custom Price Model. Besides the aforementioned common price models, DeFi applications can customize their own pricing mechanisms, which are often more diverse and complicated. For example, UwULend (UwULend, 2024) offers lending service, permitting users to borrow sUSDe by depositing cryptocurrencies as collateral. To ensure that the user can repay the debt, it is necessary to examine whether the total value of sUSDe are below that of the collateral when lending it. In this process, UwULend calculate the price of sUSDe by the following custom price model, distinct from the aforementioned CPMM and Stableswap Invariant models:
(1) $\begin{split}P_{sUSDe}=median(\{IP_{USDe,Pool_{1}},\cdots,IP_{USDe,Pool_{5}},\\ EMAP_{USDe,Pool_{1}},\cdots,EMAP_{USDe,Pool_{5}}\})\end{split}$
[…]
2.3. A Motivating Example
Figure 2. A simplified attack flow against UwULend.
Our approach is motivated by a real-world price manipulation attack on UwULend (Exvul, 2024; Etherscan, 2024a) in 2024. Figure 2 shows a attack flow of the price manipulation towards UwULend. The attack includes six key steps, and the root cause is the flawed price dependency in eq. 1 related to two functions borrow and liquidationCall.
By swapping a large amount of USDe into five liquidity pools, the attacker could deflate the instant price of USDe while the EMA prices remain unchanged, leading to a lower median price for sUSDe. Hence, the attack was able to borrow an exceedingly large amount of sUSDe. Subsequently, swapping back could increase the calculated price of sUSDe, resulting in the attacker’s collateral being unable to repay the debt , thereby allowing liquidation and acquiring deposited collateral with a bonus, leading to a \$19M loss.
While detecting a price manipulation attack, existing tools generally define the exchange rate between two different tokens as their price. For example, the price of sUSDe could be calculated by dividing the amount of WETH deposited into the protocol at step 3 by the amount of sUSDe borrowed from it at step 4. Yet, defining abnormal price change range or even precisely capturing token prices is not a trivial task.
  • DeFort (Xie et al., 2024) uses historical exchange rates, i.e., prices, between two tokens for computing a so-called normal fluctuation range. However, it may be error-prone as historical prices, especially for low-liquidity or newly created pools, could vary significantly, thus being insensitive to the detection of subtle price manipulations. For instance, in the UwULend attack, slight manipulations observed in sUSDe prices — decreasing by 4.2 % or increasing by 4.43 %, did not exceed the predefined bounds that evades the detection by DeFort illustrated in our evaluation in §7.
  • DeFiRanger (Wu et al., 2023) detects abnormal price changes by tracing token exchange sequences within a transaction and comparing token exchange rates at different steps. However, completely tracking these sequences can be challenging for complex transactions. In the UwULend attack, the attacker crafted complicated deposit and withdrawal operations concerning WETH, making it difficult for DeFiRanger to detect.
To address the aforementioned issues about price identification, our observation is that abnormal price fluctuations of tokens could be directly derived from the high-level price model and amount changes in token balances, without needing to explicitly calculate the exchange rate for a pair of tokens. Taking the UwULend attack as an example, swapping USDe into liquidity pools in step 2 decreases the value of $IP_{USDe,Pool_{i}}$, causing an abnormal drop in the median price, namely the price of sUSDe. Based on this observation, in this paper, we propose a novel price change reasoning approach powered by LLM and integrate with predefined rules to enhance the capability of price manipulation attack detection.
Figure 3. A high-level overview of DeFiScope.
3. Overview
Based on the analysis of the motivating example illustrated in §2.3, we establish an intuition of inferring price change associated with DeFi operations during the transaction process to detect DeFi price manipulations in various scenarios. While inferring through standard price models (e.g., CPMM and Stableswap as introduced in §2.2) is straightforward, it is challenging to (i) interpret low-level price calculation Solidity code into high-level price calculation formulas, and (ii) infer the price change from the (custom) price calculation formulas and the information on the token balance changes in related accounts. Both tasks require certain intelligence.
To address this key challenge, we introduce the first LLM-based approach, DeFiScope, for effective DeFi price manipulation detection. As depicted in Figure 3, DeFiScope consists of ten steps. In step ①, DeFiScope first decodes and slices raw transaction data. Then in steps ② ⑤ DeFiScope extracts the code of possible price calculation functions from smart contracts based on their signature, and token balance changes in relevant accounts. Subsequently, in steps ⑥, DeFiScope embeds these two information into a prompt template, which will be used by the LLM to extract the price model and infer the price change in steps ⑦ ⑧. In the meantime, in steps ③ ④, DeFiScope constructs the transfer graph (§5.1) and uses it to recover the high-level DeFi operations (§5.2) associated with those price change. This is because using the trend of token price changes alone is insufficient for detecting price manipulation. Finally, based on the recovered DeFi operations and their price change information, DeFiScope maps them into eight attack patterns listed in §6 and detects DeFi manipulation attacks in step ⑨.
We conduct an off-line fine-tuning step (step ⓪) to enhance LLMs’ capabilities in extracting price calculation models and token price change reasoning. This is because while the off-the-shelf LLMs may exhibit certain capability in reasoning and code understanding, they are limited for predicting the trend of token price changes given unlabeled code snippet and numerical changes of tokens balance, which will be illustrated in §7. We will detail our fine-tuning technique in §4.
4. Price Change Inference with LLMs
Instead of leveraging traditional symbolic or concrete execution techniques to estimate the price model through input and output data points, which results in a significant gap between the actual price calculation formulas and the approximating ones, we use LLMs to understand codes calculating token prices and extract corresponding price models, which eliminates the gap caused by estimation.
4.1. LLM Fine-tuning
For fine-tuning techniques, we chose OpenAI’s fine-tuning paradigm (OpenAi, 2024a) instead of supervised fine-tuning (SFT) (Chung et al., 2024) and its parameter-efficient version, LoRA (Hu et al., 2021), because a very small set of training data is required for the former while much more data points are needed for the latter (Ma et al., 2025). Moreover, OpenAI’s GPT family models demonstrate state-of-the-art reasoning capabilities on common benchmarks (Analysis, 2024; Sun et al., 2024a), which provides a suitable foundation for our fine-tuning. Accordingly, DeFiScope has enhanced the GPT-3.5-Turbo and GPT-4o models with data synthesized using the commonly used price calculation model, i.e., CPMM as illustrated in §2.2, along with on-chain data to fine-tune them.
On-Chain Data Simulation. We leverage the fuzz testing method in Foundry (Github, 2024), an off-the-shelf toolkit for Ethereum application development, to simulate on-chain data. To avoid data leakage and generate a substantial volume of transactions satisfying the CPMM, we select the Uniswap V2:BTC20 (Etherscan, 2024b) liquidity pool as our target. We randomly generate inputs, namely integers ranging from $10^{20}$ to $10^{21}$ — 100 Ether to 1000 Ether, for the swap operations which are simulated on a forked blockchain of block height 17,949,214. Specifically, to include the data of inflating the price of tokens, we craft particular operations. To begin with, we record the balance of WETH and BTC20 in the liquidity pool denoted as $bal_{WETH}$ and $bal_{BTC20}$ respectively. Then we trigger swapExactTokensForTokens in contract UniswapV2Router02 (Etherscan, 2024c) to swap a amount of BTC20 for WETH, and record the latest balance of WETH, $bal_{WETH}^{{}^{\prime}}$, and that of BTC20, $bal_{BTC20}^{{}^{\prime}}$. Finally, we obtained the tokens’ balance change as a pair ($bal_{WETH}-bal_{WETH}^{{}^{\prime}},bal_{BTC20}-bal_{BTC20}^{{}^{\prime}}$). In terms of deflating the price of tokens, we swap a amount of WETH for BTC20 instead, with similar subsequent operations. Finally, we build a fine-tuning database comprising 500 pairs for price inflating and 500 pairs for price deflating, respectively. Despite the only use of CPMM-based DeFi protocols, our evaluation results in §7 demonstrates a significant gain in term of price manipulation attack detection for DeFi protocols using custom price models.
CoT-style Fine-tuning.
Figure 4. The prompt template used in fine-tuning the LLM.
Figure 4 demonstrates the prompt template used in the fine-tuning. We construct a CoT-style fine-tuning prompt that integrates both on-chain data and the price context. Above the dashed line is the first instruction, which requires the LLM to extract the price calculation model from the provided code. {code} is the placeholder for the code snippet of price calculation functions. Below the dashed line, we guide the LLM to evaluate the credibility of four statements based on the price model extracted from step 1 and the tokens’ balance change. We demand that the LLM expresses the credibility of a statement using integers ranging from 1 to 10. Compared to merely responding with a simple “Yes” or “No,” this scoring method also indicates the confidence level of the responses, which can assist us in selecting the answers in which the LLM is more confident.
In the data part of the template, {$value_{0}$} and {$value_{1}$} are the placeholders for the first and second values in a price change pair, which are sampled from the fine-tuning dataset, respectively. {direction of change} can be either “increases” or “decreases.” Specifically, if {$value_{i}$} is greater than 0, {direction of change} is “increases”; conversely, if {$value_{i}$} is less than 0, {direction of change} is “decreases.” For the answer part, {score} is the score placeholder, which can be an integer between 1 and 10.
Following OpenAI’s fine-tuning guideline (OpenAi, 2024a) that recommends using 50 to 100 training examples, we randomly sample 96 non-repetitive (in current and previous training sets) data from the fine-tuning database, and allocate 83% of the samples for training and 17% for validation. Subsequently, we insert data from the training and validation sets into the prompt template. To obtain the desired response for each prompt, we firstly ask LLMs to generate raw response, including the analysis of price model and scores of statements, for the given prompt. Next, we manually verify the correctness of the responses, of which the correct responses are stored and the wrong responses will also be corrected. Particularly, we simply swap the scores of two opposite statements to correct the error in them. Through this process, we construct a ground truth about prompts and its responses for fine-tuning the LLM model.
During the fine-tuning process, training hyperparameters were automatically configured by the OpenAI fine-tuning API used. In particular, once the model achieved 100% accuracy on the validation set, we terminated the training to avoid overfitting. The entire fine-tuning process is cost-efficient, consuming a total of around 1 million training tokens for each model. The detailed costs across different models will be introduced in §7.2.
Figure 5. The simplified Type-I prompt and its response for the motivating example during the inference process.
4.2. Inference with the Fine-tuned LLM
In this section, we first illustrate the general inference process using the motivating example. Nevertheless, there could be closed source DeFi protocols that dissatisfy steps ⑤⑥ so that we design a customized inference process to deal with them.
General Inference Process. To assist LLMs in extracting and analyzing the price model, our inference prompt provides relevant code for LLMs to score the trend of price changes and guides them in inferring the token price changes. Figure 5 illustrates a simplified version of the prompt used and the response produced by our fine-tuned LLM for inferring price changes of the motivating example in §2.3. We call this Type-I prompt, used for the typical cases with code input retrieved from steps ⑤⑥. It is different from the fine-tuning prompt in the {statement} and {change_description} parts. In the fine-tuning prompt shown in Figure 4, both parts are fixed, while they are dynamically generated during inference using two formats: (i) “The price of {token_name} in {contract_name} {direction_of_change} after change” for the placeholder {statement}; and (ii) “The balance of {token_name} in {contract_name} {direction_of_change} by {change_value}” or “The total supply of {token_name}{direction_of_change} by {change_value}” for the placeholder {change_description}. However, to fill them in Type-I prompt, DeFiScope generates a pair of statements for each token, i.e., one regarding the increase in token price and another regarding the decrease in token price.
DeFiScope asks the fine-tuned LLM to locate the price calculation model from the input code and evaluate the credibility of the generated statements. From the motivating example’s response shown in the right-hand section of Figure 5, the LLM initially extracts the code of price calculation-related functions, followed by an high-level summary. In this example, it accurately identifies the underlying price model (see eq. 1) — the price of sUSDe is determined by the median of multiple prices. With this knowledge learned, the LLM could credit two opposite statements given the tokens balance changes and yields the correct answer with high confidence.
Figure 6. The Type-II prompt for closed-source liquidity pools.
Customized Inference Process. Although the majority of DeFi applications are open source to gain users’ trust, some remain closed source, making our Type-I prompt inapplicable. To address this, we developed a Type-II prompt template, as shown in Figure 6, to infer the trend of price changes in closed source two-token liquidity pools. Our observation is that the majority of two-token liquidity pools pools use the CPMM as their underlying price model. Therefore, the primary distinction between the Type-I prompt and the Type-II prompt lies in replacing the first instruction with a description of the liquidity pool, informing the LLM that the pool’s price model aligns with CPMM. It is worth noting that the liquidity pool is automatically identified during transaction analysis, which will be introduced in §5. Due to page limitation, we include a case study of using the Type-II prompt for inference in Appendix A.
5. DeFi Operations
The standalone fluctuations in token prices are meaningless; they need to be considered within the DeFi context to serve as evidence for detecting price manipulation. However, the raw transactions obtained from the blockchain consist solely of low-level information, such as token transfer actions and smart contract invocations. There exists a gap between raw transactions and high-level DeFi semantics.
To bridge this gap, we first model token transfer actions using a directed graph (§5.1), and then recover high-level DeFi operations from it (§5.2). Since our detection is based on a single transaction, it should be noted that all described operations are derived from one raw transaction, and we do not consider the DeFi operations expressed by the combination of multiple raw transactions. Based on our study of the top-10 high-value DeFi applications across all three categories mentioned in §2.1 (the full list is available in Appendix B) with active transactions in each category, due to the susceptibility to front-running across multiple transactions and the atomicity of transactions ensuring complete execution of operations, only a very few DeFi operations span multiple transactions.
5.1. Transfer Graph Construction
We define the Transfer Graph (TG) (Definition 5.2), a directed graph where the edges represent transfer actions (Definition 5.1) and the vertices represent related accounts, to model transfer actions within each user invocation.
Definition 0 (Transfer).
A transfer $T:=\langle s,r,t,v\rangle$, if performed successfully, deducts amount $v\in\mathbb{N}$ of token $t\in Addr$ from the sender’s account $s\in Addr$ and the balance of token $t$ in the receiver’s account $r\in Addr$ increases by $v$.
Definition 0 (Transfer Graph).
A Transfer Graph (TG) is a tuple ($\mathcal{A}$$\mathcal{E}$), where $\mathcal{A}$ is the set of all accounts (including EOAs, CAs and $\emptyset$) involved in a user invocation, $\mathcal{E}$ is the set of directed edges, i.e., =$\mathcal{E}=\{E_{1},...,E_{m}\}\subseteq\mathcal{A}\times\mathcal{A}$, where each $E_{i}:=\langle j,T_{k}\rangle$, j is the time index of $T_{k}$, $T_{k}\in\mathcal{T}$, $\mathcal{T}$ is the set of all transfer actions involved in the user invocation, i.e., $\mathcal{T}=\{T_{1},...,T_{n}\}$, where $T_{i}.s,T_{i}.r\in\mathcal{A}$ for each $T_{i}$.
According to our categorization, a transfer action can be one of three types: transferring token, burning token, and minting token. All transfer actions can be expressed as “Sender transfers amount of token to Receiver.” In a transferring token action, all accounts involved must be either EOAs or CAs, and must not be a zero address2220x0000000000000000000000000000000000000000 or a dead address3330x000000000000000000000000000000000000dEaD (we uniformly denote these two special addresses by $\emptyset$). Meanwhile, the receiver in a burning token action and the sender in a minting token action must be $\emptyset$.
Figure 7, as demonstrated in step ①, illustrates the construction of a TG from the raw transaction of a user invocation. This user invocation includes six contract accounts and a collection of user-controlled accounts $UC$, which includes EOAs and CAs, along with seven transferring token actions. The Sender and Receiver of a transfer action are connected by a directed edge, from the Sender to the Receiver, with a time index to indicate the order of occurrence. $T_{1}$, from one of the user-controlled accounts in $UC$ to $CA_{1}$, is the first transfer action in this user invocation. Before $CA_{1}$ transfers tokens to $CA_{2}$ through $T_{4}$, user-controlled accounts initiate two transfers to $CA_{4}$ and $CA_{5}$ respectively, resulting in $T_{4}$ having a larger time index compared to $T_{2}$ and $T_{3}$. Similarly, since $T_{6}$ occurs between $T_{5}$ (with a time index of 5) and $T_{7}$ (with a time index of 7), its time index is set to 6.
Compared to existing methods that model transfer actions in raw transactions utilized for price manipulation detection, such as the Cash Flow Tree (CFT) constructed in DeFiRanger (Wu et al., 2023), the novelty of the Transfer Graph (TG) lies in its unique modeling approach. The CFT models invocation and transfer actions within raw transactions, where the edges represent invocation relationships and the vertices represent a called function or a transfer action. In contrast, the TG models transfer actions in each user invocation within raw transactions, thereby offering greater granularity. Furthermore, the TG employs a directed graph rather than a tree structure for modeling and retains temporal relationships between different transfer actions, which enables it to cover more complicated DeFi operations and accelerate the speed. The proposed transfer graph are different from those fund flow graphs in prior works, such as  (McLaughlin et al., 2023). TG differs from the existing works in terms of granularity and design purpose. First, their directed graphs are built on transactions, while TG is defined on user invocations where a transaction can have multiple user invocations. Second, their directed graphs are to analyze transaction-related fund flow behaviors like arbitrage. In contrast, TG focuses on extracting high-level DeFi operations with the fund flow among user invocations.
Figure 7. An illustrative example showing the workflow of recovering DeFi operations. $UC$: a collection of user-controlled accounts, including EOAs and CAs; $CA_{i}$: the $i$-th contract account; $T_{i}$: the $i$-th transferring token action.
5.2. DeFi Operation Inference
Based on our in-depth analysis of the top-30 DeFi apps shown in Table 7 (see Appendix B), we identify six DeFi operations that need to be recovered: 1) Swap, primarily from DEXs such as Uniswap; 2) Deposit, defined by yield-farming and lending apps such as AAVE and Pendle; 3) Withdraw, mainly from yield-farming apps; 4) Borrow, based on lending protocols like Compound; 5) Stake; and 6) Claim, both predominantly from yield-farming apps like Convex Finance.
Swap involves a user exchanging one token ($Token_{in}$) for another ($Token_{out}$) from a liquidity pool. Relevant contract accounts, excluding user-controlled ones, are noted as liquidity pools for price trend analysis when source code is unavailable. In a Deposit, a user transfers a token ($Token_{deposit}$) to a yield-farming or lending protocol and receives a proof token ($Token_{proof}$) through minting. Withdraw occurs when a user retrieves tokens ($Token_{withdraw}$) from a protocol by burning a proof token ($Token_{proof}$). Borrow refers to a DeFi operation where a borrower receives a token ($Token_{borrow}$) and incurs a debt evidenced by another token ($Token_{debt}$), issued through minting. Stake happens in yield-farming protocols like Convex Finance (Etherscan, 2024a), where users deposit tokens and can earn rewards without receiving a minted proof token. Claim enables users to retrieve staked tokens ($Token_{claim}$) and bonuses without burning any tokens, unlike the Withdraw operation which requires burning tokens to retrieve assets.
An Operation Recovering Example. To recover DeFi operations from the TG, we design and employ a search algorithm based on directed graphs for each operation. Since the recovery algorithm of DeFi operations is not the central focus of this work, we use the recovery of the Swap operation as an illustrative example here.
Figure 7 illustrates that, to recover a Swap operation, we use the depth-first search algorithm to identify cycles that both start and end at user-controlled accounts. These cycles must satisfy three constraints: (i) the transfer action of each edge must be transferring token; (ii) the time index of each edge must be monotonically increasing; and (iii) the token involved in the first transfer action must differ from the token in the last transfer action. Therefore, from the well-constructed TG, we can recover the Swap operation expressed as $UC\rightarrow T_{1}\rightarrow CA_{1}\rightarrow T_{4}\rightarrow CA_{2}\rightarrow T_{5}\rightarrow CA_{3}\rightarrow T_{7}\rightarrow UC$, and we label $CA_{1}$, $CA_{2}$, and $CA_{3}$ as liquidity pools.
Table 1. Price change information-directed attack patterns, systematically mined from all types of DeFi applications (see §2.1).
Type Pattern Details
Buy & Sell  (BlockSec, 2024c) I 1) Swap $Token_{x}$ to $Token_{y}$ through $Pool_{buy}$ 2) The price of $Token_{y}$ in $Pool_{sell}$ increases / The price of $Token_{z}$ in $Pool_{sell}$ decreases 3) Swap $Token_{y}$ to $Token_{z}$ through $Pool_{sell}$
II 1) The price of $Token_{x}$ in $Pool_{buy}$ increases / The price of $Token_{y}$ in $Pool_{buy}$ decreases/       The price of $Token_{y}$ in $Pool_{sell}$ increases / The price of $Token_{z}$ in $Pool_{sell}$ decreases 2) Swap $Token_{x}$ to $Token_{y}$ through $Pool_{buy}$ 3) Swap $Token_{y}$ to $Token_{z}$ through $Pool_{sell}$
Deposit & Borrow  (Immunefi, 2024) III 1) Deposit $Token_{x}$ into $Contract_{deposit}$ and get $Token_{y}$ as credential 2) The price of $Token_{x}$ in $Contract_{borrow}$ increases / The price of $Token_{z}$ in $Contract_{borrow}$ decreases 3) Borrow $Token_{z}$ using $Token_{x}$ as collateral from $Contract_{borrow}$
IV 1) The price of $Token_{x}$ in $Contract_{borrow}$ increases / The price of $Token_{z}$ in $Contract_{borrow}$ decreases 2) Deposit $Token_{x}$ into $Contract_{deposit}$ and get $Token_{y}$ as credential 3) Borrow $Token_{z}$ using $Token_{x}$ as collateral from $Contract_{borrow}$
Stake & Claim  (Certik, 2024) V 1) Stake $Token_{x}$ into $Contract_{stake}$ 2) The price of $Token_{y}$ in $Contract_{claim}$ decreases 3) Claim $Token_{y}$ from $Contract_{claim}$
VI 1) The price of $Token_{x}$ in $Contract_{stake}$ increases / The price of $Token_{y}$ in $Contract_{claim}$ decreases 2) Stake $Token_{x}$ into $Contract_{stake}$ 3) Claim $Token_{y}$ from $Contract_{claim}$
Deposit & Withdraw  (SlowMist, 2024a) VII 1) Deposit $Token_{x}$ into $Contract_{deposit}$ and get $Token_{y}$ as credential 2) The price of $Token_{y}$ in $Contract_{withdrawal}$ increases / The price of $Token_{z}$ in $Contract_{withdrawal}$ decreases 3) Withdraw $Token_{z}$ from $Contract_{withdrawal}$ by burning $Token_{y}$
VIII 1) The price of $Token_{x}$ in $Contract_{deposit}$ increases / The price of $Token_{y}$ in $Contract_{deposit}$ decreases /       The price of $Token_{y}$ in $Contract_{withdrawal}$ increases / The price of $Token_{z}$ in $Contract_{withdrawal}$ decreases 2) Deposit $Token_{x}$ and get $Token_{y}$ as credential 3) Withdraw $Token_{z}$ from $Contract_{withdrawal}$ by burning $Token_{y}$
6. Price Manipulation Detection
The price change information and high-level DeFi operations recovered from §4 and §5, respectively, are finally checked against the detection rules (c.f. Table 1). We analyzed all four specific types of DeFi protocols (c.f. §2.1) and their associated attack instances. Specifically, we examined ElephantMoney (BlockSec, 2024c) for DEX, Cream Finance (Immunefi, 2024) for lending protocols, ATK (Certik, 2024) for staking-based yield farming protocols, and Harvest (SlowMist, 2024a) for deposit-based yield farming protocols, which collectively incurred a loss of \$163.3M.
Based on this in-depth analysis, we identify four attack types targeting different DeFi protocols and their eight generalized attack patterns, as depicted in Table 1. These attack types are Buy & Sell, Deposit & Borrow, Stake & Claim, and Deposit & Withdraw, with each type corresponding to two specific attack patterns. We will discuss these attack types and patterns in more detail.
Buy & Sell. In this type of attack strategy, the attacker primarily profits by first buying $Token_{y}$ with $Token_{x}$ through a swap in $Pool_{buy}$ and then selling $Token_{y}$ for $Token_{z}$ through a swap in $Pool_{sell}$. $Token_{x}$ and $Token_{z}$ can be the same or different tokens. In the attack against ElephantMoney, the attacker first conducted a swap in $Pool_{buy}$ to exchange WBNB for ELEPHANT and then invoked the mint function, which triggered a swap in $Pool_{sell}$ to exchange ELEPHANT for WBNB, resulting in a price increase of ELEPHANT in $Pool_{sell}$. Ultimately, the attacker utilized a reverse swap in $Pool_{sell}$ to obtain WBNB by selling ELEPHANT at the manipulated price. We design Pattern I based on this attack and subsequently generalize it to Pattern II. The major difference between these two patterns is that the token price is manipulated before the first swap in Pattern II, allowing the price of tokens in either $Pool_{buy}$ or $Pool_{sell}$ to be manipulated.
Deposit & Borrow. In this type, the attacker inflates the price of the deposited tokens or deflates the price of the borrowed assets as calculated by $Contract_{borrow}$, bypassing the protective mechanism of over-collateralization, thereby borrowing more assets than the actual value of the collateral. In the Cream Finance incident, the attacker first deposited yUSD as collateral and obtained an equivalent amount of crYUSD as proof of deposit, then inflated the price of yUSD calculated by $Contract_{borrow}$ by transferring a large quantity of yCrv to a specific contract account. Finally, using yUSD as collateral, the attacker borrowed a large amount of assets, which far exceeded the actual value of the deposited yUSD, from $Contract_{borrow}$. We design Pattern III based on this attack and then generalize it to Pattern IV. Pattern IV differs from Pattern III in that the attacker can preemptively increase the price of tokens designated for deposit or decrease the price of assets intended for borrowing as calculated by $Contract_{borrow}$ before the deposit operation. In particular, the motivating example in §2 conforms to Pattern IV.
Stake & Claim. This attack type primarily targets yield-farming protocols that offer staking services. Typically, an attacker first stakes $Token_{x}$ into the application in one transaction. The share ratio of the user is calculated based on the value and quantity of the staked asset in real-time and is stored in the state variables. Then, the attacker decreases the calculated price of $Token_{y}$ in $Contract_{claim}$ and subsequently claims $Token_{y}$ from the contract. $Token_{x}$ and $Token_{y}$ can be the same or different tokens. We derive Pattern V based on the analysis of the attack against ATK. Specifically, in the first transaction, the attacker initially staked ATK into $Contract_{stake}$. Since the staking service required that the ATK be held for 24 hours by the contract account before claiming, the attacker waited for a period and executed the second transaction, exploiting a flash loan to deflate the price of ATK in $Contract_{claim}$, subsequently claiming back an amount of ATK significantly higher than the appropriate quantity. Considering that the attacker can inflate the price of tokens intended for staking beforehand to get an incorrectly calculated share ratio, we further derive Pattern VI from Pattern V.
Deposit & Withdraw. In this attack type, the attacker exploits vulnerabilities in the token pricing mechanism within the deposit or withdrawal contract to conduct price manipulation attacks. We design and generalize Pattern VII based on the Harvest attack. In this hack, the attacker first deposited USDC ($Token_{x}$) in $Contract_{deposit}$ and received fUSDC ($Token_{y}$) as proof. Then, by exchanging USDC for USDT, the price of USDC calculated by $Contract_{withdrawal}$ decreased, and the attacker withdrew an excessive amount of USDC ($Token_{z}$) from $Contract_{withdrawal}$ by burning fUSDC. In this case, $Token_{x}$ and $Token_{z}$ are the same; however, some protocols, such as (Bscscan, 2024), allow different tokens for deposit and withdrawal. Besides deflating the price of tokens to be withdrawn, the attacker can also inflate the price of tokens used for calculating the withdrawal amount, i.e., $Token_{y}$. If the attacker manipulates the token price before depositing, the price of tokens involved in the deposit can also be affected. Based on this assumption, we generalize Pattern VII to create Pattern VIII.
7. Implementation and Evaluation
We implement DeFiScope with about 3,900 lines of Python code. DeFiScope currently supports two blockchains, i.e., Ethereum (Ethereum, 2024) and BSC (Chain, 2024), which account for over 60% of total value locked (TVL) among all the blockchains (DeFiLlama, 2024). To obtain raw data from specific transactions, we utilize blockchain node APIs facilitated by QuickNode (QuickNode, 2024), an external RPC service provider.
Our evaluation aims to answer three research questions (RQs):
  • RQ1: How effectively does DeFiScope detect price manipulation attacks compared with the existing state-of-the-arts?
  • RQ2: How significantly does the fine-tuning technique promote the accuracy of DeFiScope?
  • RQ3: How practically and efficiently does DeFiScope detect price manipulation attacks in a real-world setting?
Datasets. To address these three RQs, we collect three datasets from real DeFi transactions on Ethereum and BSC to evaluate DeFiScope, as shown in Table 2. Specifically, we use the first dataset, D1, which comprises 95 transactions of real-world price manipulation attacks from 90 DeFi applications, to evaluate both RQ1 and RQ2. To demonstrate DeFiScope’s practicality for RQ3, we use the second dataset, D2, consisting of 968 suspicious transactions collected by our industry partner. Furthermore, to measure DeFiScope’s time overhead when deployed in a realistic setting with both suspicious and benign transactions, we mix the 968 suspicious transactions from D2 with 96,800 benign transactions in the third dataset, D3.
For D1, we scraped data from multiple sources. Initially, we included all 54 price manipulation transactions identified by DeFort (Xie et al., 2024). We also collected all 55 price manipulation attacks documented by the renowned DeFi security GitHub repository, DeFiHackLabs (SunWeb3Sec, 2024), from October 26, 2020, to October 11, 2023. To further expand our dataset, we acquired 31 transactions confirmed as price manipulation attacks from our industry partner. After removing duplicates among the three sources of data, we ultimately obtained 95 price manipulation attacks for D1, which collectively caused \$381.16M in losses on Ethereum and BSC.
We also measure the distribution and monetary losses of the 95 attacks in D1 across our eight attack patterns described in §6. We observe that Pattern I has the highest number of cases, while Pattern III results in the greatest losses. This is largely due to one of the most notorious incidents, CreamFinance (Immunefi, 2024), which falls under this pattern and led to a loss of \$130M. For Patterns V and VIII, there is only one case each: ATK (Certik, 2024) and LUSD (Bscscan, 2024), respectively.
Table 2. The benchmark datasets used for evaluation.
Dataset RQs
D1: 95 historical real-world attacks RQ1, RQ2
D2: 968 suspicious transactions RQ3
D3: 96,800 benign transactions RQ3
Table 3. Summary of different attack patterns in D1.
Pattern I II III IV V VI VII VIII
#Case 49 20 5 6 1 2 11 1
Loss(\$) 55.3M 70M 141.2M 43M 61K 83K 71.4M 9K
D2 comprises 968 suspicious transactions provided by our industry partner, a Web3 security company, which monitors blockchain transactions in real-time and automatically flags transactions that yield significant profits for initiators. All these transactions are sufficiently complex and potentially involve price manipulation. While manually confirming these transactions is labor-intensive and prone to errors, 155 of them have been confirmed to belong to other vulnerabilities, such as Reentrancy (OpenLeverage (Franklin, 2024)), Unverified User Input (YIELD (BlockSec, 2024e)), and Access Control Bugs (SafeMoon (zokyo, 2024)), by developers and industry partners. In this way, we can stress-test DeFiScope’s detection capabilities when mixed with other types of attack transactions.
Furthermore, to assess DeFiScope’s false alarms on benign transactions and to measure its time overhead in a realistic setting, we need to collect a large number of benign transactions, as the most majority of real-world transactions are still benign. In the absence of empirical data on the ratio of suspicious to benign transactions, we adopt a conservative yet reasonable ratio of 1:100, assuming one suspicious transaction for every 100 benign transactions. Thus, with 968 suspicious transactions in D2, we require 96,800 benign transactions for D3. To this end, we randomly sampled these 96,800 transactions from DeFort (Xie et al., 2024)’s dataset of 428,523 benign transactions, which includes 384,143 benign transactions on Ethereum and BSC. At this sample size, we achieve a 99.999% confidence level with a margin of error of 0.625%.
Experimental Setup. All experiments were conducted on a desktop computer running Ubuntu 20.04, powered by an Intel® Xeon® W-2235 CPU (3.80 GHz, 6 cores, and 12 threads) and equipped with 16 GB of memory. For LLMs, we use OpenAI’s GPT family models due to limited data required for fine-tuning, as explained in §4.1. DeFiScope by default uses GPT-3.5-Turbo (GPT-3.5-turbo-1106) for its well-recognized cost-performance balance, but we also conduct an ablation study in §7.2 to test the more advanced GPT-4o (GPT-4o-2024-08-06) for fine-tuning. For the LLM configuration, DeFiScope employs nucleus sampling (Holtzman et al., 2019) with a top-p value of 1 and sets the temperature to 0, ensuring highly deterministic responses for each prompt, although each was run only once.
Table 4. Detection results for 95 ground-truth DeFi price manipulation attacks. DS: DeFiScope, DT: DeFiTainter, DF: DeFort, and DR: DeFiRanger. DeFiScope detected 76 attacks, DeFiTainter detected 34 attacks, DeFort detected 50 attacks, and DeFiRanger detected 49 attacks. Note that DeFiScope failed in 8 cases solely due to ① missing source code and ② compilation errors, which were not caused by DeFiScope’s methodology.
Protocol Chain Date Loss DS DT DF DR
AES BSC 07-Dec-22 60K
APC BSC 01-Dec-22 6K
APEDAO BSC 18-Jul-23 7K
ApeRocket BSC 14-Jul-21 1.26M
ARK BSC 23-Mar-24 201K
ArrayFinance ETH 18-Jul-21 516K
ATK BSC 12-Oct-22 61K
AutoSharkFinance_1 BSC 29-Oct-21 2M
AutoSharkFinance_2 BSC 24-May-21 750K
BabyDoge BSC 28-May-23 137K
Bamboo BSC 04-Jul-23 117K
BBOX BSC 16-Nov-22 12K
BDEX BSC 30-Oct-22 3K
bDollar BSC 21-May-22 730K
BEARNDAO BSC 05-Dec-23 769K
BeltFinance_1 BSC 29-May-21 408K
BeltFinance_2 BSC 29-May-21 6.23M
BFCToken BSC 09-Sep-23 38K
BGLD BSC 12-Dec-22 18K
BH BSC 11-Oct-23 1.27M
BTC20 ETH 19-Aug-23 47K
BXH BSC 28-Sep-22 40K
bZx ETH 18-Feb-20 350K
Carson BSC 26-Jul-23 150K
Cellframe BSC 01-Jun-23 76K
CheeseBank ETH 06-Nov-20 3.3M
ConicFinance ETH 21-Jul-23 3.25M
CreamFinance ETH 27-Oct-21 130M
CS BSC 23-May-23 714K
Cupid BSC 31-Aug-22 78K
DFS BSC 30-Dec-22 2K
Discover BSC 06-Jun-22 11K
DotFinance BSC 25-Aug-21 430K
EAC BSC 29-Aug-23 17K
EGDFinance BSC 07-Aug-22 36K
ElephantMoney BSC 12-Apr-22 11.2M
Eminence ETH 29-Sep-20 7M
ERC20TokenBank ETH 31-May-23  [^1] K
FFIST BSC 19-Jul-23 91K
GDS BSC 03-Jan-23 180K
GPT BSC 24-May-23 42K
Groker20 ETH 10-Nov-23 68K
GymDeFi BSC 09-Apr-22 312K
Hackerdao BSC 24-May-22 65K
Harvest ETH 26-Oct-20 21.5M
IndexedFinance ETH 14-Oct-21 16M
INUKO BSC 14-Oct-22 50K
InverseFinance_1 ETH 16-Jun-22 1.26M
InverseFinance_2 ETH 02-Apr-22 15.6M
Protocol Chain Date Loss DS DT DF DR
LaunchZone BSC 27-Feb-23 320K
LUSD BSC 07-Jul-23 9K
LW_1 BSC 12-May-23 50K
LW_2 BSC 12-May-23 48K
Mars BSC 16-Apr-24 100K
MBC BSC 29-Nov-22 6K
MerlinLab BSC 29-Jun-21 628K
MonoXFinance ETH 30-Nov-21 31M
MRGtoken ETH 08-Nov-23 12K
NeverFall BSC 02-May-23 74K
Nmbplatform BSC 14-Dec-22 76K
NOVO_1 BSC 29-May-22 76K
NOVO_2 BSC 29-May-22 65K
PancakeBunny BSC 19-May-21 45M
PancakeHunny BSC 20-Oct-21 1.93M
PLPManager BSC 24-Jul-23 900K
PLTD BSC 17-Oct-22 24K
RoeFinance ETH 11-Jan-23 80K
SanshuInu ETH 20-Jul-21  [^1] K
SATX BSC 16-Apr-24 29K
SellToken BSC 11-Jun-23 100K
SpaceGodzilla BSC 13-Jul-22 25K
SpartanProtocol BSC 01-May-21 30M
Starlink BSC 16-Feb-23 12K
StarWallets BSC 17-Apr-24 33K
STM BSC 06-Jun-24 14K
SturdyFinance ETH 12-Jun-23 800K
SVT BSC 26-Aug-23 400K
SwapX BSC 27-Feb-23 1M
TIFIToken BSC 10-Dec-22 51K
UN BSC 06-Jun-23 26K
UPSToken ETH 18-Jan-23 45K
Upswing ETH 17-Jan-23 36K
uwerx_network ETH 02-Aug-23 324K
UwULend ETH 10-Jun-24 19M
ValueDeFi ETH 14-Nov-20 6M
VesperFinance ETH 02-Nov-21 2M
VINU ETH 06-Jun-23 6K
WarpFinance ETH 17-Dec-20 7.8M
WGPT BSC 12-Jul-23 80K
WienerDoge BSC 25-Apr-22 30K
XSTABLE ETH 09-Aug-22 56K
Z123 BSC 22-Apr-24 136K
Zoompro BSC 05-Sep-22 61K
ZS BSC 08-Oct-23 14K
Zunami ETH 13-Aug-23 2M
7.1. RQ1: Detection Effectiveness
To answer RQ1, we evaluate DeFiScope and compare it with three SOTA tools—DeFiTainter (Kong et al., 2023), DeFort (Xie et al., 2024), and DeFiRanger (Wu et al., 2023), using the dataset D1. Because DeFiRanger is not open-source, we use the results reported in their paper for the attacks they evaluated; for other attacks, we re-implemented their approach based on the description (Wu et al., 2023) and conducted our evaluation. Although DeFort is also not open-source, we obtained a copy of its source code from the authors to conduct our experiment.
Table 4 presents the detection results. The first four columns list the name of the victim protocol, the destination chain, the hack date, and the resulting loss, respectively. Note that one protocol may face multiple attacks, so we add numerical suffixes to protocol names to differentiate them. We use ✓ to indicate an attack can be successfully detected by a tool, and ✗ to indicate a detection failure.
DeFiScope can detect most of the price manipulation attacks. It achieves a recall rate of 80%, outperforming all other tools. Overall, DeFiScope detected 76 attacks, followed by DeFort with 50 and DeFiRanger with 49 attacks, respectively, while DeFiTainter detected only 34 attacks. Figure 8 details the performance of each tool on the evaluated DeFi protocols across four application categories using different price models. Compared to other tools, DeFiScope performs the best in every application category. Particularly, it achieves the highest recall, 90.7%, in Token-related protocols. However, DeFiScope yields a low recall rate in Lending-related protocols, though still higher than all existing tools. Through further analysis, we identifies that 3 out of 12 attacks (InverseFinance_2, SanshuInu, and VesperFinance) are cross-transaction attacks that DeFiScope is unable to detect. Additionally, one attack (TIFIToken) involves exploiting a closed-source custom price model, and the InverseFinance_1 attack, a false negative, will be detailed below.
Yet, DeFiScope missed detecting 11 price manipulation attacks. After analyzing each attack, we discovered that 8 out of 11 are cross-transaction attacks (Certik, 2024; Ancilia, 2024b; Phalcon, 2024; Inspex, 2024; Knownsec, 2024; Ancilia, 2024c; BlockSec, 2024d; Finance, 2024), where detection was unsuccessful because DeFiScope is based on analyzing individual transactions. For Zoompro, detection failure occurred because the token involved in the transaction did not adhere to the ERC20 token standard, resulting in the transfer event not being identified. The remaining two attacks, i.e., “IndexedFinance” and “InverseFinance_1” cases, were subject to in-depth analysis. The detection incapability for IndexedFinance was due to its use of an extremely complicated pricing mechanism that involves exponential calculations in its price-related function joinswapExternAmountIn. In the case of InverseFinance, the attacker exploited the flawed price dependency when calculating the price of tokens deposited as collateral, where the collateral price is based on the balance of multiple tokens in the liquidity pool and off-chain price oracles. As both cases require precise quantitative calculations, DeFiScope is limited by the current LLMs’ constrained capacity for scientific computation. A potential solution could be to integrate with Program-Aided Language models (PAL) (Gao et al., 2023), guiding the LLM to generate scripts for necessary calculations and executing them to obtain the result.
Among the remaining 8 undetectable transactions, 3 were not analyzed successfully due to the unavailability of the code of price calculation functions since our method relies on code-level analysis; five cases involved compilation errors during the extraction of price calculation functions using Slither (crytic, 2024), an off-the-shelf static analyzer for Solidity and Vyper, thereby prematurely terminating the detection process. These limitations are not inherent to DeFiScope’s methodology and can be mitigated by automated or semi-automated techniques, e.g., code decompilation (Grech et al., 2019, 2022; Suiche, 2017) and manual intervention.
Figure 8. The categorized detection results on DeFi protocols.
7.2. RQ2: Ablation Study
In this RQ, we investigate how fine-tuning can enhance DeFiScope’s detection accuracy on the same D1 dataset, as well as the impact and cost of fine-tuning different GPT models. To this end, we test four settings shown in Figure 9: (a) the original GPT-3.5-Turbo without fine-tuning, (b) GPT-3.5-Turbo with fine-tuning, which was used in RQ1, (c) the original GPT-4o without fine-tuning, and (d) GPT-4o with fine-tuning.
Figure 9 shows that fine-tuning significantly enhances DeFiScope’s ability to detect attacks, with the fine-tuned versions of GPT-3.5-Turbo and GPT-4o detecting 18 (31%) and 12 (19%) more attacks, respectively. It also indicates that fine-tuning provides a more noticeable improvement for less powerful models, such as GPT-3.5-Turbo, compared to stronger models like GPT-4o. With fine-tuning, the detection success rate for attacks targeting the CPMM increases to 100% with GPT-3.5-Turbo and 95.6% with GPT-4o [^4] GPT-4o does not exhibit an advantage with CPMM but shows a clear advantage with custom price models, especially when the original GPT-4o compared to GPT-3.5-Turbo.. This very high rate can be attributed to the use of CPMM during the fine-tuning training phase, enabling the model to effectively handle this specific pricing model.
(a) GPT-3.5-Turbo w/o fine-tuning
(b) GPT-3.5-Turbo w/ fine-tuning
(c) GPT-4o w/o fine-tuning
(d) GPT-4o w/ fine-tuning
Figure 9. Effectiveness of fine-tuning for 78 attacks that go through DeFiScope’s LLM inference. The cases not plotted in the figures include eight due to missing source code and compilation errors, and nine due to cross-transaction issues and non-adherence to the ERC20 token standard, as mentioned in §7.1.
Although DeFiScope’s fine-tuned models were trained using only CPMM data, we find that they also exhibit strong transfer learning capabilities for attacks targeting custom price models. Specifically, the detection success rate for attacks targeting custom price models increases from 60% to 93.3% (a relative increase of 55.5%) for GPT-3.5-Turbo and from 86.7% to 96.7% (a relative increase of 11.5%) for GPT-4o. Further analysis of the LLM responses generated during detection reveals that the most significant difference introduced by fine-tuning is that the fine-tuned model strictly adheres to the CoT approach we specified. This involves initially extracting the price model from the given code and then conducting inference based on the extracted model along with the provided information about balance changes. In contrast, although the original model produces the final evaluation scores, it does not strictly follow the instructions of the CoT prompts.
Table 5. Comparison between GPT-3.5-Turbo and GPT-4o.
Model $TP$ $Recall$ Ave. Cost(\$)/ per inference Fine-tuning Cost(\$)
GPT-3.5 w/o fine-tuning 58 0.61 0.0023 -
GPT-3.5 w/ fine-tuning 76 0.80 0.0107 8
GPT-4o w/o fine-tuning 63 0.66 0.008 -
GPT-4o w/ fine-tuning 75 0.79 0.0131 25
LLM Costs. Table 5 highlights that while the original GPT-4o generally performs better than GPT-3.5-Turbo, this performance gap can be largely minimized through fine-tuning. However, regarding costs, GPT-3.5-Turbo has a clear advantage, which is why DeFiScope uses GPT-3.5-Turbo for fine-tuning by default instead of GPT-4o. Specifically, the fine-tuning cost of GPT-3.5-Turbo is 68% lower than that of GPT-4o, and the per-request inference cost of GPT-3.5-Turbo is merely \$0.0107 (OpenAi, 2024b), which is also 18% cheaper than that of GPT-4o. Therefore, for a tradeoff between cost and performance, we recommend using GPT-3.5-Turbo for DeFiScope.
7.3. RQ3: Real-World Practicality
To be practical in a real-world setting, DeFiScope not only needs to maintain high detection rates for true attacks but also minimize false alarms for benign transactions. In this RQ, we evaluate this aspect of DeFiScope and the associated time overhead using the datasets D2 and D3, which were introduced in the prologue of §7. Specifically, D2 comprises 968 suspicious transactions with various attacks (i.e., not limited to price manipulations), and D3 includes 96,800 benign transactions collected from DeFort (Xie et al., 2024)’s dataset.
For D2, DeFiScope flagged 153 price manipulation attacks out of 968 suspicious transactions. To robustly confirm these potential attacks, we cross-referenced them with attack reports or alerts published by security companies through their official channels (Ancilia, 2024a; Beosin, 2024; BlockSec, 2024b; SlowMist, 2024b), thereby verifying the root causes of the attacks. Using this method, we confirm that 66 of them are previously reported price manipulation attacks. For the remaining cases, we conducted comprehensive and in-depth analyses by combining manual review with ancillary evidence, such as identifying whether the EOA initiating the transaction was marked as a hacker by blockchain explorers (BscScan, 2024; EtherScan, 2024; OKLink, 2024). Finally, we discovered 81 previously unknown historical incidents, which were neither reported by security companies nor tagged as malicious transactions by blockchain explorers, and identified a total of six false positive cases (as detailed in Table 6 in Appendix C). Of these, five are benign transactions initiated by the same EOA and exhibiting identical invocation flows, while the last one is indeed an attack due to a logic issue rather than a price manipulation attack, resulting in a precision of 96% (147/153).
In comparison, DeFort identified only 58 attacks, of which 9 were false positives labeled as other vulnerabilities by security researchers. Moreover, DeFort failed to detect 114 of the malicious transactions flagged by DeFiScope, including 54 that are officially confirmed attacks.
We further analyze DeFiScope’s false alarms in a realistic setting mixing 968 suspicious transactions in D2 with 96,800 benign transactions in D3 (see the earlier dataset setup in the prologue of this section). For each transaction, we set the maximum scan time to 300 seconds. The results reveal that DeFiScope achieves zero false alarms on benign transactions in D3, with an average of 2.5 seconds per transaction across all transactions. This highlights DeFiScope’s potential for large-scale, daily on-chain monitoring scenarios.
8. Discussion
While the precision of DeFiScope is quite satisfactory, its recall rate primarily affected by the following two threats. Future work can further optimize the following two aspects:
Cross-transaction price manipulation attacks. DeFiScope currently detects only single-transaction attacks. However, some price manipulation attacks are designed across multiple transactions to circumvent time restrictions coded in DeFi protocols, as studied in (Chen et al., 2024b). Detecting cross-transaction attacks is often challenging. For example, attacks like INUKO (Ancilia, 2024b) span 48 hours (crossing around 57,000 blocks), making it extremely difficult to precisely identify all related transactions. Fortunately, our study observed that cross-transaction attacks are less common than single-transaction ones.
Closed-source price calculation functions. The availability of code for price models could affect the detection accuracy of DeFiScope. According to our study, most DeFi applications are open-source to gain user trust. To mitigate issues with closed-source price models, we design the Type-II prompt to cover price models in those closed-source liquidity pools.
9. Related Work
On-chain Security Analysis. Prior efforts (Zhou et al., 2021b; Qin et al., 2021, 2022; Daian et al., 2020; Zhou et al., 2023) have elucidated threats in blockchain and attacks against the DeFi ecosystem. Some transaction-based systems have been proposed to mine vulnerable transaction sequences (Zhang et al., 2020; So et al., 2021), explore arbitrage opportunities (Zhou et al., 2021a), detect malicious phishing (He et al., 2023), and simulate attacks to prevent intrusions (Qin et al., 2023a).
Tools like FlashSyn (Chen et al., 2024a), DeFiRanger (Wu et al., 2023), DeFiTainter (Kong et al., 2023), and DeFort (Xie et al., 2024) are capable of detecting attacks associated with price manipulation. FlashSyn uses numerical approximation techniques to synthesize malicious contracts that target DeFi apps through price manipulation attacks. However, there is a gap between the synthesized contracts and those that launch attacks on the live blockchain.DeFiRanger generates cash flow trees (CFT) from a transaction, translates the semantics of CFT into high-level DeFi actions, and uses specific patterns to detect price manipulation attacks. Yet, DeFiRanger is significantly limited by its reliance on predefined templates. DeFiTainter constructs call graphs based on contract states and transactions, then performs inter-contract taint analysis to examine all execution paths and uncover price manipulation attacks. Nonetheless, DeFiTainter’s approach requires manual analysis and labeling of taint sources and sinks, which is challenging. DeFort uses a price manipulation behavior model to guide detection and analysis but its boundary calculation mechanism lacks sensitivity to attacks characterized by minor price fluctuations.
Program Analysis for Contract Security. Traditional tools such as Slither (Feist et al., 2019), Zeus (Kalra et al., 2018), and Securify (Tsankov et al., 2018) employ static analysis methods to detect vulnerabilities at the source code and bytecode levels. Symbolic execution has also shown significant potential in this area, with tools like Mythril (Consensys, 2024), Oyente (Luu et al., 2016), and Manticore (Mossberg et al., 2019) developed based on this technique. Dynamic analysis is another core methodology utilized by various tools (Jiang et al., 2018; Liu et al., 2018; Shou et al., 2023; Qin et al., 2023b). Some tools are specifically designed to detect particular types of vulnerabilities; for example, VetSC (Duan et al., 2022) identifies business logic bugs by vetting inconsistencies between UI and contract code, SailFish (Bose et al., 2022) focuses on state inconsistency issues, and JACKAL (Gritti et al., 2023) targets confused deputy vulnerabilities.
LLMs for Smart Contract Security. LLMs have become formidable tools in the realm of blockchain security. GPTScan (Sun et al., 2024b) uses LLMs for code understanding, incorporating static analysis to detect logical vulnerabilities in contracts. BlockGPT (Gai et al., 2023) serves as a real-time intrusion detection system, using LLMs to rank transaction anomalies. LLM4Vuln (Sun et al., 2024a) refines the reasoning processes of LLMs when analyzing smart contract vulnerabilities. iAudit (Ma et al., 2025) combines fine-tuning with LLM-driven agents to provide intuitive smart contract auditing with explanations. PropertyGPT (Liu et al., 2025) uses retrieval-augmented property generation for smart contracts, leveraging LLMs’ in-context learning capabilities. Similarly, SMARTINV (Wang et al., 2024) employs Tier of Thought (ToT) to reason across different modalities of smart contracts, generating invariants in the process. To the best of our knowledge, DeFiScope is the first tool that uses LLMs specifically designed for detecting price manipulation attacks.
10. Conclusion
In this paper, we introduced DeFiScope, the first tool that utilizes LLMs specifically for detecting price manipulation attacks. DeFiScope employs LLMs to intelligently infer the trend of token price changes based on balance information within transaction executions. To strengthen LLMs in this aspect, we simulated on-chain transaction data and fine-tuned a DeFi price-specific LLM. We also proposed a graph-based method to recover high-level DeFi operations and systematically mined eight price manipulation patterns. Our evaluation demonstrated DeFiScope’s superior performance over SOTA approaches and real-world impact. Future work includes better handling of closed-source price calculation functions.
Appendix
Table 6. Details of six false positives mentioned in §7.3.
Transaction hash Chain Block Type Root Cause
0x130c6370 BSC 38218540 Benign -
0x4b59af93 BSC 38218538 Benign -
0xe158a2b9 BSC 38218537 Benign -
0x59942848 BSC 38218536 Benign -
0x2e9ceb16 BSC 38218539 Benign -
0x640ce34c BSC 11403670 Malicious Logic issue
Appendix A A Case Study of Type-II Prompt
Figure 10 illustrates an use case of Type-II prompt. The contract (0x2120…3379) is closed source DEX contract that allows users to trade SVT and USDT. While recovering the DeFi operations, this contract was identified as a two-token liquidity pool. So, we applied the Type-II prompt to inferring the token price trend in its transactions. As shown by the LLM response in Figure 10, the primary difference from the response to a Type-I prompt lies in the absence of the analysis on the price model; instead, the scoring is directly yielded. The is because the prompt itself assumes that the CPMM is employed in this contract. Regarding the CPMM, the relationship between token price and token balance is quite standardized, i.e., the direction of balance change is opposite to that of price change.
Figure 10. An example of Type-II prompt and its response generated while detecting the attack against SVT (BlockSec, 2024a).
Table 7. High-value DeFi applications (from DefiLlama (DefiLlama, 2024)) with active transactions of each category we have done a comprehensive study on. TVL refers to the Total Value Locked in the app.
DEX App TVL Lending App TVL Yield farming App TVL
Uniswap \$4.8B Compound \$2.0B Pendle \$2.8B
Curve \$1.9B AAVE \$12.3B Convex Finance \$1.1B
Balancer \$777.2M Morpho \$1.5B Aura \$361.9M
Sushi \$250.2M Fraxlend \$134.4M Magpie \$192.5M
PancakeSwap \$1.7B Venus \$1.5B StakeDAO \$78.8M
1inch \$4.58M Strike \$9.7M Equilibria Finance \$80.0M
ParaSwap \$6.34M Planet \$1.2M Kine Finance \$8.0M
ShibaSwap \$18.19M Kinza Finance \$42.8M Dot Dot Finance \$2.1M
BiSwap \$27.1M Radiant \$7.8M Solo Top \$1.8M
MDEX \$16.0M Ambit \$6M Jetfuel Finance \$1.5M
Appendix B High-value DeFi Applications
The design of DeFi operations and the category of transfer actions in §5 is based on an in-depth study of the high-value DeFi applications listed in Table 7 as of August 2024.
Appendix C Details of False Positives
Table 6 presents the details of six false positives discovered in our experiments described in §7.3. The reason for these false positives is that their transactions involve two contract accounts that were created three months ago by the transaction initiator. Any fund transfers among these accounts and the initiator should be considered benign operations rather than price manipulation operations. Yet, these accounts were incorrectly marked as closed-source DEXes in the detection, leading to false inferences. Such false positives could be mitigated by conducting a historical analysis of account ownership relationships and clustering user accounts that are controlled by the same owners.

### v2

@@
-DeFiScope: Detecting Various DeFi Price Manipulations with LLM Reasoning
-Abstract.
+Detecting Various DeFi Price Manipulations with LLM Reasoning
+Abstract
 DeFi (Decentralized Finance) is one of the most important applications of today’s cryptocurrencies and smart contracts.
@@
 In this paper, we introduce the first LLM-based approach, DeFiScope, for detecting DeFi price manipulation attacks in both standard and custom price models.
-Our insight is that large language models (LLMs) have certain intelligence to abstract price calculation from code and infer the trend of token price changes based on the extracted price models.
+Our insight is that large language models (LLMs) have certain intelligence to abstract price calculation from smart contract source code and infer the trend of token price changes based on the extracted price models.
 To further strengthen LLMs in this aspect, we leverage Foundry to synthesize on-chain data and use it to fine-tune a DeFi price-specific LLM.
 Together with the high-level DeFi operations recovered from low-level transaction data, DeFiScope detects various DeFi price manipulations according to systematically mined patterns.
-Experimental results show that DeFiScope achieves a high precision of 96% and a recall rate of 80%, significantly outperforming SOTA approaches.
+Experimental results show that DeFiScope achieves a high recall of 80% on real-world attacks, a precision of 96% on suspicious transactions, and zero false alarms on benign transactions, significantly outperforming SOTA approaches.
 Moreover, we evaluate DeFiScope’s cost-effectiveness and demonstrate its practicality by helping our industry partner confirm 147 real-world price manipulation attacks, including discovering 81 previously unknown historical incidents.
-copyright: none
-1. Introduction
-DeFi represents a form of finance that eliminates traditional financial intermediaries by utilizing smart contracts on a blockchain.
-A smart contract (Zou et al., 2019) is a self-executing program with the terms of the agreement between the buyer and the seller directly written into lines of code.
-The code and the agreements contained therein exist across a distributed, decentralized blockchain network (Dinh et al., 2018).
-Smart contracts enable trusted transactions and agreements to be carried out among disparate, anonymous parties without the need for a central authority, legal system, or external enforcement mechanism.
-While price manipulation is a pervasive risk in traditional financial markets, it poses even more significant damage in the context of DeFi.
-In the cryptocurrency space, attackers strategically add or remove large amounts of tokens in specific contracts, or impact the total supply of target tokens by burning or minting, to exploit vulnerabilities in the pricing mechanisms of DeFi applications, ultimately inflating or deflating the price of cryptocurrencies (Cointelegraph, 2024).
-To detect price manipulation attacks in DeFi applications, researchers have proposed several state-of-the-art (SOTA) approaches, primarily focusing on transaction monitoring-based methods that offer real-time protection, e.g., DeFiRanger (Wu et al., 2023) and DeFort (Xie et al., 2024).
+I Introduction
+DeFi represents a form of finance that eliminates traditional intermediaries by utilizing smart contracts on a blockchain.
+A smart contract [1] is a self-executing program with the terms of the agreement directly written into code, deployed on a distributed, decentralized blockchain network [2].
+These contracts power decentralized applications such as Automated Market Makers (AMMs), where token prices are determined by on-chain pricing mechanisms rather than centralized order books.
+While price manipulation is a pervasive risk in traditional finance, DeFi pricing mechanisms create more attack surfaces: adversaries can strategically add or remove liquidity, or impact token supply by minting or burning, to exploit vulnerabilities in pricing mechanisms, ultimately distorting prices of cryptocurrencies [3].
+To detect price manipulation attacks in DeFi applications, researchers have proposed several state-of-the-art (SOTA) approaches, primarily focusing on transaction monitoring-based methods that offer real-time protection, e.g., DeFiRanger [4] and DeFort [5].
 However, our analysis indicates that these SOTA approaches are less effective to non-standard price models in custom DeFi protocols.
-This is because they typically require token exchange rates to identify abnormal price changes, which are only suitable to calculate under standard price models such as CPMM (Constant Product Market Makers) and Stableswap Invariant (detailed in §2.2).
+This is because they typically require token exchange rates to identify abnormal price changes, which are only suitable to calculate under standard price models such as CPMM (Constant Product Market Makers) and Stableswap Invariant (detailed in §II-A).
 Unfortunately, our subsequent evaluation shows that 44.2% of the 95 DeFi price manipulation attacks reported in the past three years used non-standard price models.
@@
 To achieve this objective, we sought help from large language models (LLMs), considering that their trained intelligence might aid in inferring price changes associated with DeFi operations during the transaction process.
-As such, we introduce DeFiScope, the first LLM-based approach to detecting DeFi price manipulation attacks in both standard and custom price models.
-DeFiScope features several novel designs, such as constructing a transfer graph to recover high-level DeFi operations (§5) and systematically mining price manipulation patterns across all types of DeFi applications (§6).
-Among them, the key design is to fine-tune a DeFi price-specific LLM (§4), where we propose (i) simulating transactions using Foundry (Github, 2024) to generate our own fine-tuning data, and (ii) conducting a Chain-of-Thought (CoT)-style fine-tuning that integrates both on-chain data and the price context.
+As such, we introduce DeFiScope, a source code-level detector and the first LLM-based approach to detecting DeFi price manipulation attacks in both standard and custom price models.
+DeFiScope features several novel designs, such as constructing a transfer graph to recover high-level DeFi operations (§V) and systematically mining price manipulation patterns across four prevalent types of DeFi applications (§VI).
+Among them, the key design is to fine-tune a DeFi price-specific LLM (§IV), where we propose (i) simulating transactions using Foundry [6] to generate our own fine-tuning data, and (ii) conducting a Chain-of-Thought (CoT)-style fine-tuning that integrates both on-chain data and the price context.
 The ablation study shows that fine-tuning increases the overall detection effectiveness by up to 31% compared to the baseline LLM (under DeFiScope) without fine-tuning.
@@
 In the benchmark experiment, we collected 95 real-world price manipulation attacks from multiple sources and compared DeFiScope with three SOTA tools.
-The results show that DeFiScope achieves a superior detection accuracy (recall) of 80%, significantly higher than 51.6% with DeFiRanger, 52.6% with DeFort, and 35.8% with DeFiTainter (Kong et al., 2023) (a static source code analysis tool).
-In a large-scale experiment involving 968 suspicious transactions (a mix of various DeFi attacks, not just price manipulations) from our industry partner and 96,800 benign transactions from DeFort (Xie et al., 2024), we helped confirm 147 out of 153 price manipulation attacks reported by DeFiScope, yielding a precision of 96%.
-In particular, 81 of these attacks are previously unknown historical incidents.
-Moreover, DeFiScope achieves zero false alarms on benign transactions, with an average of 2.5-second time overhead across suspicious and benign transactions and merely \$0.0107 per-request LLM inference cost.
+The results show that DeFiScope achieves a superior detection accuracy (recall) of 80%, significantly higher than 51.6% with DeFiRanger, 52.6% with DeFort, and 35.8% with DeFiTainter [7] (a static source code analysis tool).
+In our large-scale experiment, we evaluated DeFiScope on two datasets provided by our industry partner and DeFort [5].
+The first dataset consists of 968 suspicious transactions, covering a variety of DeFi attacks beyond price manipulations.
+On this dataset, DeFiScope flagged 153 potential price manipulation attacks, 147 of which were manually verified as true positives, yielding a precision of 96%.
+Notably, 81 of these attacks are previously undocumented price manipulation incidents.
+The second dataset involves 96,800 benign transactions, on which DeFiScope produced zero false alarms.
+Regarding efficiency, DeFiScope incurs an average time overhead of 2.5 seconds per transaction across both suspicious and benign transactions, with a per-request LLM inference cost of only \$0.0107.
 In sum, this paper makes the following contributions:
@@
   • We extensively evaluate DeFiScope with three real-world transaction datasets, showing DeFiScope’s superior performance over SOTA tools in terms of precision and recall.
-Open Science and Ethics.
-To facilitate future research, all source code and experimental results are available in an (anonymous) GitHub repository (Anonymous, 2024).
-For ethical research, we ensure that all attacks elaborated in this work have been published or investigated by authoritative security companies through their established channels, and did not involve any undisclosed zero-day vulnerabilities.
-2. Background and Motivation
-As Figure 1 shows, on blockchains, two primary account types exist, i.e., externally owned accounts (EOAs), controlled by individuals using private keys, and contract accounts (CAs), governed by their contract code.
+Artifact. To facilitate open science and future research, all source code, experimental results, and supplementary material are available in https://github.com/AIS2Lab/DeFiScope.
+II Background and Motivation
+On blockchains, two primary account types exist, i.e., externally owned accounts (EOAs), controlled by individuals using private keys, and contract accounts (CAs), governed by their contract code.
 External transactions are initiated externally by EOAs, while internal transactions occurs when a smart contract calls another contract internally.
-Smart contracts exchange information with other smart contracts through internal transactions, where a transaction involving communication between contracts can be regarded as a sequence of function calls, which has been greatly utilized in DeFi protocols for interoperability.
+Smart contracts exchange information with other smart contracts through internal transactions, where a transaction involving communication between contracts can be regarded as a sequence of function calls, utilized by DeFi protocols for interoperability.
 For simplicity, we denote as user invocation an internal transaction between the user-controlled smart contracts and other smart contracts, which plays a key role in price manipulation attacks.
-2.1. DeFi Protocols
-A DeFi protocol typically consists of multiple smart contracts and offers various services including trading on decentralized exchange, lending, and yield farming.
-Decentralized EXchange (DEX).
-A DEX is a peer-to-peer marketplace where users can trade cryptocurrencies.
-The two predominant forms of DEXs are order book DEXs and automated market makers (AMMs).
-In the order book DEXs, the management and matching processes of open buy and sell orders take place off-chain while the settlement of trades occurs on-chain.
-Alternatively, the AMM is fully on-chain, and maintains liquidity through liquidity pools without replying on real-time buyers and sellers.
-A liquidity pool maintains two or more tokens, and the prices of these tokens are determined by a constant mathematical formula.
-In this paper, we only target fully on-chain DEXs using AMM (Chainlink, 2024; coinbase, 2024).
-Lending.
-To borrow assets from a lending protocol, borrowers are required to deposit other cryptocurrencies as collateral first.
-To ensure that borrowers can repay their loans, the collateral factor of supplied assets is utilized to calculate the maximum value of assets that users can borrow.
-For instance, consider a scenario where the collateral factor of ETH is set to 0.75, borrowers can only obtain assets up to 75% of the value of ETH they have deposited as collateral.
-If the value of the collateral decreases and falls below a certain level (the health factor of a borrower is less than 1), the liquidation of the collateral (the lending protocol sells the collateral at a discount rate) occurs to repay the borrowed amount (moai finance, 2024).
-Yield Farming.
-It provides a way to earn rewards in the form of Annual Percentage Yields (APYs).
-A yield farmer could deposit cryptocurrencies into yield-farming applications, which invest these assets according to various strategies.
-Based on the returns from these investments, yield farmers can obtain rewards.
-There are two types of yield farming.
-The first involves users depositing cryptocurrencies into an application to obtain proof tokens, which can be traded, allowing holders to withdraw the corresponding deposited assets.
-The second, a stake-and-claim model, has users lock their assets into the contract and later claim them back.
-This process does not generate proof tokens, meaning the staked shares are non-tradable, and only the staking account can make the claim.
-Figure 1. The blockchain transactions.
-2.2. Price Models
+II-A Price Models
 A price model represents the pricing mechanism within the DeFi application, which are typically expressed as equations correlating the price of a certain token with the balances of and the total supply of various tokens as well as other constant.
 Constant Product Market Maker (CPMM).
-CPMM is one of the most prevalent DeFi AMMs and used in many well-known DEXs (Uniswap, 2024; PancakeSwap, 2024).
-It maintains liquidity through a constant function $R_{x}*R_{y}=k$, where $k$ is a constant, and $R_{x}$ and $R_{y}$ represent the reserves of two distinct assets $token_{x}$ and $token_{y}$ in a liquidity pool.
-In a swap operation, let $\Delta x$ amount of $token_{x}$ can exchange for $\Delta y$ amount of $token_{y}$, resulting in $(x+\Delta x)*(y-\Delta y)=k$ where the instantaneous price of $token_{x}$ denominated in $token_{y}$ is $P_{x,y}=\frac{y}{x}$ (Wu et al., 2023; Gogol et al., 2024).
-Stableswap Invariant. The Stableswap Invariant (Curve.fi, 2024c), widely used in Curve AMMs (Curve.fi, 2024a), and is defined as follows:
+CPMM is one of the most prevalent DeFi AMMs and used in many well-known DEXs [8, 9].
+It maintains liquidity through a constant function xy=kx*y=k, where $k$ is a constant, and $x$ and $y$ represent the reserves of two distinct assets $token_{x}$ and $token_{y}$ in a liquidity pool.
+In a swap operation, let $\Delta x$ amount of $token_{x}$ can exchange for $\Delta y$ amount of $token_{y}$, resulting in $(x+\Delta x)*(y-\Delta y)=k$ where the instantaneous price of $token_{x}$ denominated in $token_{y}$ is $P_{x,y}=\frac{y}{x}$ [4, 10].
+Stableswap Invariant. The Stableswap Invariant [11], widely used in Curve AMMs [12], and is defined as follows:
 $\frac{a\prod_{i=1}^{n}x_{i}}{(D/n)^{n}}\cdot D^{n-1}\cdot\sum_{i=1}^{n}x_{i}+\prod_{i=1}^{n}x_{i}$ $=\frac{a\prod_{i=1}^{n}x_{i}}{(D/n)^{n}}\cdot D^{n}+\frac{D^{n}}{n^{n}}$
-where $a$ indicates a constant amplification coefficient and n is the number of tokens in the liquidity pool, while $D$ represents the total underlying value, i.e., the total amount of tokens in the pool when token prices are equal, and $x_{i}$ denotes the current reserve of $token_{i}$.
-As $a$ approaches zero, the function behaves similarly to a CPMM.
+where $a$ indicates a constant amplification coefficient and $n$ is the number of tokens in the liquidity pool, while $D$ represents the total amount of tokens in the pool when token prices are equal, and $x_{i}$ denotes the current reserve of $token_{i}$.
+The curve is nearly flat in the middle part, while price changes at the extremes resemble CPMM behavior.
 Custom Price Model.
-Besides the aforementioned common price models, DeFi applications can customize their own pricing mechanisms, which are often more diverse and complicated.
-For example, UwULend (UwULend, 2024) offers lending service, permitting users to borrow sUSDe by depositing cryptocurrencies as collateral.
-To ensure that the user can repay the debt, it is necessary to examine whether the total value of sUSDe are below that of the collateral when lending it.
-In this process, UwULend calculate the price of sUSDe by the following custom price model, distinct from the aforementioned CPMM and Stableswap Invariant models:
-(1) $\begin{split}P_{sUSDe}=median(\{IP_{USDe,Pool_{1}},\cdots,IP_{USDe,Pool_{5}},\\ EMAP_{USDe,Pool_{1}},\cdots,EMAP_{USDe,Pool_{5}}\})\end{split}$
+Besides the mentioned common price models, DeFi protocols can customize their own pricing mechanisms, which are often more diverse and complicated.
+For example, UwULend [13] offers lending service, permitting users to borrow sUSDe by depositing cryptocurrencies as collateral.
+To ensure the user can repay, examining whether the total value of sUSDe are below that of the collateral when lending it is necessary.
+In this process, UwULend calculates the price of sUSDe by the following custom price model, distinct from the aforementioned CPMM and Stableswap Invariant models:
+$\begin{split}P_{sUSDe}=median(\{IP_{USDe,Pool_{1}},\cdots,IP_{USDe,Pool_{5}},\\ EMAP_{USDe,Pool_{1}},\cdots,EMAP_{USDe,Pool_{5}}\})\end{split}$ (1)
 […]
-2.3. A Motivating Example
-Figure 2. A simplified attack flow against UwULend.
-Our approach is motivated by a real-world price manipulation attack on UwULend (Exvul, 2024; Etherscan, 2024a) in 2024.
-Figure 2 shows a attack flow of the price manipulation towards UwULend.
-The attack includes six key steps, and the root cause is the flawed price dependency in eq.
+II-B A Motivating Example
+Our approach is motivated by a real-world price manipulation attack on UwULend [20, 21] in 2024.
+The root cause of the attack is the flawed price dependency in eq.
 1 related to two functions borrow and liquidationCall.
-By swapping a large amount of USDe into five liquidity pools, the attacker could deflate the instant price of USDe while the EMA prices remain unchanged, leading to a lower median price for sUSDe.
-Hence, the attack was able to borrow an exceedingly large amount of sUSDe.
-Subsequently, swapping back could increase the calculated price of sUSDe, resulting in the attacker’s collateral being unable to repay the debt , thereby allowing liquidation and acquiring deposited collateral with a bonus, leading to a \$19M loss.
+By swapping a large amount of USDe into five liquidity pools, the attacker deflated the instant price of USDe while the EMA prices kept unchanged, resulting in a lower median price for sUSDe.
+Hence, the attack was able to borrow an exceedingly large amount of sUSDe using the same amount of collateral.
+Then, swapping back could increase the calculated price of sUSDe, resulting in the attacker’s collateral being unable to repay the debt, thereby allowing liquidation and acquiring deposited collateral with a bonus, leading to a \$19M loss.
 While detecting a price manipulation attack, existing tools generally define the exchange rate between two different tokens as their price.
-For example, the price of sUSDe could be calculated by dividing the amount of WETH deposited into the protocol at step 3 by the amount of sUSDe borrowed from it at step 4.
+For example, the price of sUSDe could be calculated by dividing the amount of WETH deposited into the protocol by the amount of sUSDe borrowed from it.
 Yet, defining abnormal price change range or even precisely capturing token prices is not a trivial task.
-  • DeFort (Xie et al., 2024) uses historical exchange rates, i.e., prices, between two tokens for computing a so-called normal fluctuation range.
+Figure 1: A high-level overview of DeFiScope.
+  • DeFort [5] uses historical exchange rates, i.e., prices, between two tokens to compute a so-called normal fluctuation range.
 However, it may be error-prone as historical prices, especially for low-liquidity or newly created pools, could vary significantly, thus being insensitive to the detection of subtle price manipulations.
-For instance, in the UwULend attack, slight manipulations observed in sUSDe prices — decreasing by 4.2 % or increasing by 4.43 %, did not exceed the predefined bounds that evades the detection by DeFort illustrated in our evaluation in §7.
-  • DeFiRanger (Wu et al., 2023) detects abnormal price changes by tracing token exchange sequences within a transaction and comparing token exchange rates at different steps.
-However, completely tracking these sequences can be challenging for complex transactions.
+For instance, in the UwULend attack, slight manipulations observed in sUSDe prices — decreasing by 4.2 % or increasing by 4.43 %, did not exceed the predefined bounds that evades DeFort’s detection, illustrated in our evaluation in §VII.
+  • DeFiRanger [4] detects abnormal price changes by tracing token exchange sequences within a transaction and comparing token exchange rates.
+However, completely tracing these sequences is challenging for complex transactions.
 In the UwULend attack, the attacker crafted complicated deposit and withdrawal operations concerning WETH, making it difficult for DeFiRanger to detect.
 To address the aforementioned issues about price identification, our observation is that abnormal price fluctuations of tokens could be directly derived from the high-level price model and amount changes in token balances, without needing to explicitly calculate the exchange rate for a pair of tokens.
-Taking the UwULend attack as an example, swapping USDe into liquidity pools in step 2 decreases the value of $IP_{USDe,Pool_{i}}$, causing an abnormal drop in the median price, namely the price of sUSDe.
+Taking the UwULend attack as an example, swapping USDe into liquidity pools decreases the value of $IP_{USDe,Pool_{i}}$, causing an abnormal drop in the median price, namely the price of sUSDe.
 Based on this observation, in this paper, we propose a novel price change reasoning approach powered by LLM and integrate with predefined rules to enhance the capability of price manipulation attack detection.
-Figure 3. A high-level overview of DeFiScope.
-3. Overview
-Based on the analysis of the motivating example illustrated in §2.3, we establish an intuition of inferring price change associated with DeFi operations during the transaction process to detect DeFi price manipulations in various scenarios.
-While inferring through standard price models (e.g., CPMM and Stableswap as introduced in §2.2) is straightforward, it is challenging to (i) interpret low-level price calculation Solidity code into high-level price calculation formulas, and (ii) infer the price change from the (custom) price calculation formulas and the information on the token balance changes in related accounts.
+III Overview
+Based on the analysis of the motivating example illustrated in §II-B, we establish an intuition of inferring price change associated with DeFi operations during the transaction process to detect DeFi price manipulations in various scenarios.
+While inferring through standard price models (i.e., CPMM and Stableswap as introduced in §II-A) is straightforward, it is challenging to (i) interpret low-level price calculation Solidity code into high-level price calculation formulas, and (ii) infer the price change from the (custom) price calculation formulas and the information on the token balance changes in related accounts.
 Both tasks require certain intelligence.
 To address this key challenge, we introduce the first LLM-based approach, DeFiScope, for effective DeFi price manipulation detection.
-As depicted in Figure 3, DeFiScope consists of ten steps.
-In step ①, DeFiScope first decodes and slices raw transaction data.
+As depicted in Fig.
+1, DeFiScope consists of ten steps.
+In step ①, DeFiScope first decodes and slices raw transaction data, and retrieves the source code of smart contracts called within the transaction.
 Then in steps ② ⑤ DeFiScope extracts the code of possible price calculation functions from smart contracts based on their signature, and token balance changes in relevant accounts.
 Subsequently, in steps ⑥, DeFiScope embeds these two information into a prompt template, which will be used by the LLM to extract the price model and infer the price change in steps ⑦ ⑧.
-In the meantime, in steps ③ ④, DeFiScope constructs the transfer graph (§5.1) and uses it to recover the high-level DeFi operations (§5.2) associated with those price change.
+In the meantime, in steps ③ ④, DeFiScope constructs the transfer graph (§V-A) and uses it to recover the high-level DeFi operations (§V-B) associated with those price change.
 This is because using the trend of token price changes alone is insufficient for detecting price manipulation.
-Finally, based on the recovered DeFi operations and their price change information, DeFiScope maps them into eight attack patterns listed in §6 and detects DeFi manipulation attacks in step ⑨.
+Finally, based on the recovered DeFi operations and their price change information, DeFiScope maps them into eight attack patterns listed in §VI and detects DeFi manipulation attacks in step ⑨.
 We conduct an off-line fine-tuning step (step ⓪) to enhance LLMs’ capabilities in extracting price calculation models and token price change reasoning.
-This is because while the off-the-shelf LLMs may exhibit certain capability in reasoning and code understanding, they are limited for predicting the trend of token price changes given unlabeled code snippet and numerical changes of tokens balance, which will be illustrated in §7.
-We will detail our fine-tuning technique in §4.
-4. Price Change Inference with LLMs
-Instead of leveraging traditional symbolic or concrete execution techniques to estimate the price model through input and output data points, which results in a significant gap between the actual price calculation formulas and the approximating ones, we use LLMs to understand codes calculating token prices and extract corresponding price models, which eliminates the gap caused by estimation.
-4.1. LLM Fine-tuning
-For fine-tuning techniques, we chose OpenAI’s fine-tuning paradigm (OpenAi, 2024a) instead of supervised fine-tuning (SFT) (Chung et al., 2024) and its parameter-efficient version, LoRA (Hu et al., 2021), because a very small set of training data is required for the former while much more data points are needed for the latter (Ma et al., 2025).
-Moreover, OpenAI’s GPT family models demonstrate state-of-the-art reasoning capabilities on common benchmarks (Analysis, 2024; Sun et al., 2024a), which provides a suitable foundation for our fine-tuning.
-Accordingly, DeFiScope has enhanced the GPT-3.5-Turbo and GPT-4o models with data synthesized using the commonly used price calculation model, i.e., CPMM as illustrated in §2.2, along with on-chain data to fine-tune them.
+This is because while the off-the-shelf LLMs may exhibit certain capability in reasoning and code understanding, they are limited for predicting the trend of token price changes given unlabeled code snippet and numerical changes of tokens balance, which will be illustrated in §VII-B.
+We now detail our fine-tuning technique in §IV.
+IV Price Change Inference with LLMs
+IV-A LLM Fine-tuning
+For fine-tuning techniques, we by default use OpenAI’s fine-tuning paradigm [22] instead of supervised fine-tuning (SFT) [23, 24], because the former requires only a small set of training data, whereas the latter generally requires more data points [25].
+Nonetheless, we also demonstrate the generalizability of DeFiScope’s fine-tuning with open-source SFT models; see details in §VIII.
+In the default setting, DeFiScope enhances the GPT-3.5-Turbo and GPT-4o models with data synthesized using the CPMM price model as illustrated in §II-A, along with on-chain data to fine-tune them.
+Our fine-tuning process comprises two components: (a) on-chain data simulation, which generates realistic token balance-change pairs from transactions on a forked blockchain network; and (b) chain-of-thought (CoT)-style fine-tuning, where these simulated data are integrated into CoT-style prompts that are then used to fine-tune the LLM, enabling it to perform stepwise reasoning.
 On-Chain Data Simulation.
-We leverage the fuzz testing method in Foundry (Github, 2024), an off-the-shelf toolkit for Ethereum application development, to simulate on-chain data.
-To avoid data leakage and generate a substantial volume of transactions satisfying the CPMM, we select the Uniswap V2:BTC20 (Etherscan, 2024b) liquidity pool as our target.
+We leverage the fuzz testing method in Foundry [6], an off-the-shelf toolkit for Ethereum application development, to simulate on-chain data.
+To avoid data leakage and generate a substantial volume of transactions satisfying the CPMM, we select the Uniswap V2:BTC20 [26] liquidity pool as our target.
 We randomly generate inputs, namely integers ranging from $10^{20}$ to $10^{21}$ — 100 Ether to 1000 Ether, for the swap operations which are simulated on a forked blockchain of block height 17,949,214.
@@
 To begin with, we record the balance of WETH and BTC20 in the liquidity pool denoted as $bal_{WETH}$ and $bal_{BTC20}$ respectively.
-Then we trigger swapExactTokensForTokens in contract UniswapV2Router02 (Etherscan, 2024c) to swap a amount of BTC20 for WETH, and record the latest balance of WETH, $bal_{WETH}^{{}^{\prime}}$, and that of BTC20, $bal_{BTC20}^{{}^{\prime}}$.
+Then we trigger swapExactTokensForTokens in contract UniswapV2Router02 [27] to swap a amount of BTC20 for WETH, and record the latest balance of WETH, $bal_{WETH}^{{}^{\prime}}$, and that of BTC20, $bal_{BTC20}^{{}^{\prime}}$.
 Finally, we obtained the tokens’ balance change as a pair ($bal_{WETH}-bal_{WETH}^{{}^{\prime}},bal_{BTC20}-bal_{BTC20}^{{}^{\prime}}$).
 In terms of deflating the price of tokens, we swap a amount of WETH for BTC20 instead, with similar subsequent operations.
-Finally, we build a fine-tuning database comprising 500 pairs for price inflating and 500 pairs for price deflating, respectively.
-Despite the only use of CPMM-based DeFi protocols, our evaluation results in §7 demonstrates a significant gain in term of price manipulation attack detection for DeFi protocols using custom price models.
+Finally, we build a fine-tuning database with 500 pairs each for price inflating and deflating.
+Despite the only use of CPMM-based DeFi protocols, our evaluation results in §VII demonstrates a significant gain in term of price manipulation attack detection for DeFi protocols using custom price models.
+Figure 2: The simplified prompt template used in fine-tuning the LLM. The full version can be found in our supplementary material [28].
 CoT-style Fine-tuning.
-Figure 4. The prompt template used in fine-tuning the LLM.
-Figure 4 demonstrates the prompt template used in the fine-tuning.
-We construct a CoT-style fine-tuning prompt that integrates both on-chain data and the price context.
+Fig.
+2 demonstrates the prompt template used in the fine-tuning.
+We carefully construct a CoT-style fine-tuning prompt for guiding the LLM toward step-by-step reasoning, which integrates both simulated on-chain data collected in the previous step and the price context.
 Above the dashed line is the first instruction, which requires the LLM to extract the price calculation model from the provided code.
@@
 We demand that the LLM expresses the credibility of a statement using integers ranging from 1 to 10.
-Compared to merely responding with a simple “Yes” or “No,” this scoring method also indicates the confidence level of the responses, which can assist us in selecting the answers in which the LLM is more confident.
+Compared to merely responding with a simple “Yes” or “No”, this scoring method also indicates the confidence level of the responses, which can assist us in selecting the answers in which the LLM is more confident.
+By offering a finer-grained evaluation scale, it reveals when the model is uncertain—if two contradictory statements receive the same confidence score, we know it cannot distinguish which is more trustworthy and therefore discard those low-confidence results.
 In the data part of the template, $value_{0}$} and $value_{1}$} are the placeholders for the first and second values in a price change pair, which are sampled from the fine-tuning dataset, respectively.
-{direction of change} can be either “increases” or “decreases.” Specifically, if {$value_{i}$} is greater than 0, {direction of change} is “increases”; conversely, if {$value_{i}$} is less than 0, {direction of change} is “decreases.” For the answer part, {score} is the score placeholder, which can be an integer between 1 and 10.
-Following OpenAI’s fine-tuning guideline (OpenAi, 2024a) that recommends using 50 to 100 training examples, we randomly sample 96 non-repetitive (in current and previous training sets) data from the fine-tuning database, and allocate 83% of the samples for training and 17% for validation.
+{direction of change} can be either “increases” or “decreases.” Specifically, if $value_{i}$} is greater than 0, {direction of change} is “increases”; conversely, if $value_{i}$} is less than 0, {direction of change} is “decreases.” For the answer part, {score} is the score placeholder, an integer between 1 and 10.
+Following OpenAI’s fine-tuning guideline [22] that recommends using 50 to 100 training examples, we randomly sample 96 non-repetitive (in current and previous training sets) data from the fine-tuning database, and allocate 83% of the samples for training and 17% for validation.
 Subsequently, we insert data from the training and validation sets into the prompt template.
@@
 Particularly, we simply swap the scores of two opposite statements to correct the error in them.
-Through this process, we construct a ground truth about prompts and its responses for fine-tuning the LLM model.
-During the fine-tuning process, training hyperparameters were automatically configured by the OpenAI fine-tuning API used.
+Through this process, we construct a ground truth about prompts and its responses for fine-tuning.
+During the fine-tuning process, training hyperparameters were automatically configured by the OpenAI fine-tuning API (for SFT fine-tuning, hyperparameters were manually set using common practice; see §VIII).
 In particular, once the model achieved 100% accuracy on the validation set, we terminated the training to avoid overfitting.
 The entire fine-tuning process is cost-efficient, consuming a total of around 1 million training tokens for each model.
-The detailed costs across different models will be introduced in §7.2.
-Figure 5. The simplified Type-I prompt and its response for the motivating example during the inference process.
-4.2. Inference with the Fine-tuned LLM
-In this section, we first illustrate the general inference process using the motivating example.
+The detailed costs across different models will be introduced in §VII-B.
+Figure 3: The simplified Type-I prompt and its response for the motivating example during the inference process.
+IV-B Inference with the Fine-tuned LLM
+We first illustrate the general inference process using the motivating example.
 Nevertheless, there could be closed source DeFi protocols that dissatisfy steps ⑤⑥ so that we design a customized inference process to deal with them.
@@
 To assist LLMs in extracting and analyzing the price model, our inference prompt provides relevant code for LLMs to score the trend of price changes and guides them in inferring the token price changes.
-Figure 5 illustrates a simplified version of the prompt used and the response produced by our fine-tuned LLM for inferring price changes of the motivating example in §2.3.
+Fig.
+3 illustrates a simplified version of the prompt used and the response produced by our fine-tuned LLM for inferring price changes of the motivating example in §II-B.
 We call this Type-I prompt, used for the typical cases with code input retrieved from steps ⑤⑥.
-It is different from the fine-tuning prompt in the {statement} and {change_description} parts.
-In the fine-tuning prompt shown in Figure 4, both parts are fixed, while they are dynamically generated during inference using two formats: (i) “The price of {token_name} in {contract_name} {direction_of_change} after change” for the placeholder {statement}; and (ii) “The balance of {token_name} in {contract_name} {direction_of_change} by {change_value}” or “The total supply of {token_name}{direction_of_change} by {change_value}” for the placeholder {change_description}.
-However, to fill them in Type-I prompt, DeFiScope generates a pair of statements for each token, i.e., one regarding the increase in token price and another regarding the decrease in token price.
+It is different from the fine-tuning prompt in the <statement> and <change_description> parts.
+In the fine-tuning prompt shown in Fig.
+2, both parts are fixed, while they are dynamically generated during inference using two formats: (i) “The price of {token_name} in {contract_name} {direction_of_change} after change” for the part <statement>, and (ii) “The balance of {token_name} in {contract_name} {direction_of_change} by {change_value}” or “The total supply of {token_name} {direction_of_change} by {change_value}” for the part <change_description>.
+Specifically, to fill them in Type-I prompt, DeFiScope generates a pair of statements for each token, i.e., one regarding the increase in token price and another regarding the decrease in token price.
 DeFiScope asks the fine-tuned LLM to locate the price calculation model from the input code and evaluate the credibility of the generated statements.
-From the motivating example’s response shown in the right-hand section of Figure 5, the LLM initially extracts the code of price calculation-related functions, followed by an high-level summary.
-In this example, it accurately identifies the underlying price model (see eq.
-1) — the price of sUSDe is determined by the median of multiple prices.
-With this knowledge learned, the LLM could credit two opposite statements given the tokens balance changes and yields the correct answer with high confidence.
-Figure 6. The Type-II prompt for closed-source liquidity pools.
+From the motivating example’s response shown in the right-hand section of Fig.
+3, the LLM first extracts the code of price calculation-related functions, followed by a high-level summary.
+In this example, it accurately identifies the underlying price model (see Equation 1): the price of sUSDe is determined by the median of multiple prices.
+With this knowledge, the LLM can evaluate two opposing statements based on the changes in token balances and provide the correct answer with high confidence, i.e., a higher score.
 Customized Inference Process.
 Although the majority of DeFi applications are open source to gain users’ trust, some remain closed source, making our Type-I prompt inapplicable.
-To address this, we developed a Type-II prompt template, as shown in Figure 6, to infer the trend of price changes in closed source two-token liquidity pools.
-Our observation is that the majority of two-token liquidity pools pools use the CPMM as their underlying price model.
-Therefore, the primary distinction between the Type-I prompt and the Type-II prompt lies in replacing the first instruction with a description of the liquidity pool, informing the LLM that the pool’s price model aligns with CPMM.
-It is worth noting that the liquidity pool is automatically identified during transaction analysis, which will be introduced in §5.
-Due to page limitation, we include a case study of using the Type-II prompt for inference in Appendix A.
-5. DeFi Operations
+To address this, we developed a Type-II prompt template to infer the trend of price changes in closed source two-token liquidity pools.
+Our observation is that the majority of two-token liquidity pools use the CPMM as their underlying price model.
+Therefore, the primary distinction between the Type-I prompt and the Type-II prompt lies in replacing the Instruction 1 with a description of the liquidity pool, informing the LLM that the pool’s price model aligns with CPMM.
+It is worth noting that the liquidity pool is automatically identified during transaction analysis, which will be introduced in §V.
+To evaluate its effectiveness, we applied Type-II on a test dataset [29] of 100 price changes on CPMM, and it reached 97% accuracy, close to the 99% accuracy of Type-I.
+This demonstrates that, if applied to previously identified two-token liquidity pools, which commonly adopt CPMM, Type-II is substantially effective comparable to Type-I, while also adapting to closed-source pools.
+More details about the Type-II prompt and a case study of using it are available in [28].
+V DeFi Operations
 The standalone fluctuations in token prices are meaningless; they need to be considered within the DeFi context to serve as evidence for detecting price manipulation.
@@
 There exists a gap between raw transactions and high-level DeFi semantics.
-To bridge this gap, we first model token transfer actions using a directed graph (§5.1), and then recover high-level DeFi operations from it (§5.2).
+To bridge this gap, we first model token transfer actions using a directed graph (§V-A), and then recover high-level DeFi operations from it (§V-B).
 Since our detection is based on a single transaction, it should be noted that all described operations are derived from one raw transaction, and we do not consider the DeFi operations expressed by the combination of multiple raw transactions.
-Based on our study of the top-10 high-value DeFi applications across all three categories mentioned in §2.1 (the full list is available in Appendix B) with active transactions in each category, due to the susceptibility to front-running across multiple transactions and the atomicity of transactions ensuring complete execution of operations, only a very few DeFi operations span multiple transactions.
-5.1. Transfer Graph Construction
-We define the Transfer Graph (TG) (Definition 5.2), a directed graph where the edges represent transfer actions (Definition 5.1) and the vertices represent related accounts, to model transfer actions within each user invocation.
-Definition 0 (Transfer).
+Based on our study of the top-10 high-value DeFi applications (the full list could be found in [28]) across three categories — Decentralized EXchange (DEX), Lending, Yield Farming, with active transactions in each category, due to the susceptibility to front-running across multiple transactions and the atomicity of transactions ensuring complete execution of operations, only a very few DeFi operations span multiple transactions.
+V-A Transfer Graph Construction
+We define the Transfer Graph (TG) (Definition 2), a directed graph where the edges represent transfer actions (Definition 1) and the vertices represent related accounts, to model transfer actions within each user invocation.
+Definition 1 (Transfer)
 A transfer $T:=\langle s,r,t,v\rangle$, if performed successfully, deducts amount $v\in\mathbb{N}$ of token $t\in Addr$ from the sender’s account $s\in Addr$ and the balance of token $t$ in the receiver’s account $r\in Addr$ increases by $v$.
-Definition 0 (Transfer Graph).
-A Transfer Graph (TG) is a tuple ($\mathcal{A}$$\mathcal{E}$), where $\mathcal{A}$ is the set of all accounts (including EOAs, CAs and $\emptyset$) involved in a user invocation, $\mathcal{E}$ is the set of directed edges, i.e., =$\mathcal{E}=\{E_{1},...,E_{m}\}\subseteq\mathcal{A}\times\mathcal{A}$, where each $E_{i}:=\langle j,T_{k}\rangle$, j is the time index of $T_{k}$, $T_{k}\in\mathcal{T}$, $\mathcal{T}$ is the set of all transfer actions involved in the user invocation, i.e., $\mathcal{T}=\{T_{1},...,T_{n}\}$, where $T_{i}.s,T_{i}.r\in\mathcal{A}$ for each $T_{i}$.
-According to our categorization, a transfer action can be one of three types: transferring token, burning token, and minting token.
-All transfer actions can be expressed as “Sender transfers amount of token to Receiver.” In a transferring token action, all accounts involved must be either EOAs or CAs, and must not be a zero address2220x0000000000000000000000000000000000000000 or a dead address3330x000000000000000000000000000000000000dEaD (we uniformly denote these two special addresses by $\emptyset$).
-Meanwhile, the receiver in a burning token action and the sender in a minting token action must be $\emptyset$.
-Figure 7, as demonstrated in step ①, illustrates the construction of a TG from the raw transaction of a user invocation.
+Definition 2 (Transfer Graph)
+A Transfer Graph (TG) is a tuple ($\mathcal{A}$$\mathcal{E}$), where $\mathcal{A}$ is the set of all accounts (including EOAs, CAs and $\emptyset$) involved in a user invocation, $\mathcal{E}$ is the set of directed edges, i.e., $\mathcal{E}=\{E_{1},...,E_{m}\}\subseteq\mathcal{A}\times\mathcal{A}$, where each $E_{i}:=\langle j,T_{k}\rangle$, j is the time index of $T_{k}$, $_{k}\in\mathcal{T}, 𝒯\mathcal{T}$ is the set of all transfer actions involved in the user invocation, i.e., $\mathcal{T}=\{T_{1},...,T_{n}\}$, where $T_{i}.s,T_{i}.r\in\mathcal{A}$ for each $T_{i}$.
+According to our categorization, a transfer action can be one of three types: transferring, burning, and minting token.
+All transfer actions can be expressed as “Sender transfers amount of token to Receiver.” In the transferring token, all accounts involved must be either EOAs or CAs, and must not be a zero address or a dead address (we uniformly denote these two special addresses by $\emptyset$).
+Meanwhile, the receiver in burning token and the sender in minting token actions must be $\emptyset$.
+Fig.
+4, as demonstrated in step ①, illustrates the construction of a TG from the raw transaction of a user invocation.
 This user invocation includes six contract accounts and a collection of user-controlled accounts $UC$, which includes EOAs and CAs, along with seven transferring token actions.
@@
 Similarly, since $T_{6}$ occurs between $T_{5}$ (with a time index of 5) and $T_{7}$ (with a time index of 7), its time index is set to 6.
-Compared to existing methods that model transfer actions in raw transactions utilized for price manipulation detection, such as the Cash Flow Tree (CFT) constructed in DeFiRanger (Wu et al., 2023), the novelty of the Transfer Graph (TG) lies in its unique modeling approach.
-The CFT models invocation and transfer actions within raw transactions, where the edges represent invocation relationships and the vertices represent a called function or a transfer action.
-In contrast, the TG models transfer actions in each user invocation within raw transactions, thereby offering greater granularity.
-Furthermore, the TG employs a directed graph rather than a tree structure for modeling and retains temporal relationships between different transfer actions, which enables it to cover more complicated DeFi operations and accelerate the speed.
-The proposed transfer graph are different from those fund flow graphs in prior works, such as  (McLaughlin et al., 2023).
-TG differs from the existing works in terms of granularity and design purpose.
-First, their directed graphs are built on transactions, while TG is defined on user invocations where a transaction can have multiple user invocations.
-Second, their directed graphs are to analyze transaction-related fund flow behaviors like arbitrage.
-In contrast, TG focuses on extracting high-level DeFi operations with the fund flow among user invocations.
-Figure 7.
-An illustrative example showing the workflow of recovering DeFi operations.
-$UC$: a collection of user-controlled accounts, including EOAs and CAs; $CA_{i}$: the $i$-th contract account; $T_{i}$: the $i$-th transferring token action.
-5.2. DeFi Operation Inference
-Based on our in-depth analysis of the top-30 DeFi apps shown in Table 7 (see Appendix B), we identify six DeFi operations that need to be recovered: 1) Swap, primarily from DEXs such as Uniswap; 2) Deposit, defined by yield-farming and lending apps such as AAVE and Pendle; 3) Withdraw, mainly from yield-farming apps; 4) Borrow, based on lending protocols like Compound; 5) Stake; and 6) Claim, both predominantly from yield-farming apps like Convex Finance.
+Compared to the CFT approach used in DeFiRanger [4], which models invocation and transfer actions within raw transactions, our TG provides finer granularity.
+TG represents transfers as a directed graph while preserving their temporal order, enabling it to capture more complex operations more effectively.
+In our evaluation on 1,000 real-world Ethereum transactions, TG achieved a TPR ($\frac{\#TP}{\#TP+\#FN}$) of 0.912, significantly higher than CFT’s 0.559, while maintaining comparable precision ($\frac{\#TP}{\#TP+\#FP}$), i.e., 0.984 vs.
+0.991.
+TG further demonstrated substantial advantages in recovering complex swap operations involving multiple relayers, achieving an 89.3% improvement over CFT.
+More details about the experimental setup and analysis are available in  [28].
+Compared to prior fund-flow graph approaches, such as  [30], TG differs in two key aspects: (i) it is constructed at the level of user invocations rather than transactions, which may contain multiple invocations, and (ii) it is designed for high-level DeFi operation recovery, whereas those primarily target behavior analysis (e.g., arbitrage).
+Figure 4: An illustrative example showing the workflow of recovering DeFi operations.
+UCUC: a collection of user-controlled accounts, including EOAs and CAs; $CA_{i}$: the $i$-th contract account; $T_{i}$: the $i$-th transferring token action.
+V-B DeFi Operation Inference
+TABLE I: Price change information-directed attack patterns.
+Type Pattern Details
+Buy
+Sell
+ [31]
+I
+1) Swap $Token_{x}$ to $Token_{y}$ through $Pool_{buy}$
+2) The price of $Token_{y}$ in $Pool_{sell}$ increases / The price of $Token_{z}$ in $Pool_{sell}$ decreases
+3) Swap $Token_{y}$ to $Token_{z}$ through $Pool_{sell}$
+II
+1) The price of $Token_{x}$ in $Pool_{buy}$ increases / The price of $Token_{y}$ in $Pool_{buy}$ decreases/
+   The price of $Token_{y}$ in $Pool_{sell}$ increases / The price of $Token_{z}$ in $Pool_{sell}$ decreases
+2) Swap $Token_{x}$ to $Token_{y}$ through $Pool_{buy}$
+3) Swap $Token_{y}$ to $Token_{z}$ through $Pool_{sell}$
+Deposit
+Borrow
+ [32]
+III
+1) Deposit $Token_{x}$ into $Contract_{deposit}$ and get $Token_{y}$ as credential
+2) The price of $Token_{x}$ in $Contract_{borrow}$ increases / The price of $Token_{z}$ in $Contract_{borrow}$ decreases
+3) Borrow $Token_{z}$ using $Token_{x}$ as collateral from $Contract_{borrow}$
+IV
+1) The price of $Token_{x}$ in $Contract_{borrow}$ increases / The price of $Token_{z}$ in $Contract_{borrow}$ decreases
+2) Deposit $Token_{x}$ into $Contract_{deposit}$ and get $Token_{y}$ as credential
+3) Borrow $Token_{z}$ using $Token_{x}$ as collateral from $Contract_{borrow}$
+Stake
+Claim
+ [33]
+V
+1) Stake $Token_{x}$ into $Contract_{stake}$
+2) The price of $Token_{y}$ in $Contract_{claim}$ decreases
+3) Claim $Token_{y}$ from $Contract_{claim}$
+VI
+1) The price of $Token_{x}$ in $Contract_{stake}$ increases / The price of $Token_{y}$ in $Contract_{claim}$ decreases
+2) Stake $Token_{x}$ into $Contract_{stake}$
+3) Claim $Token_{y}$ from $Contract_{claim}$
+Deposit
+Withdraw
+ [34]
+VII
+1) Deposit $Token_{x}$ into $Contract_{deposit}$ and get $Token_{y}$ as credential
+2) The price of $Token_{y}$ in $Contract_{withdrawal}$ increases / The price of $Token_{z}$ in $Contract_{withdrawal}$ decreases
+3) Withdraw $Token_{z}$ from $Contract_{withdrawal}$ by burning $Token_{y}$
+VIII
+1) The price of $Token_{x}$ in $Contract_{deposit}$ increases / The price of $Token_{y}$ in $Contract_{deposit}$ decreases /
+   The price of $Token_{y}$ in $Contract_{withdrawal}$ increases / The price of $Token_{z}$ in $Contract_{withdrawal}$ decreases
+2) Deposit $Token_{x}$ and get $Token_{y}$ as credential
+3) Withdraw $Token_{z}$ from $Contract_{withdrawal}$ by burning $Token_{y}$
+Based on our in-depth analysis of the top-30 DeFi apps, we identify six DeFi operations that need to be recovered: 1) Swap, primarily from DEXs such as Uniswap; 2) Deposit, defined by yield-farming and lending apps such as AAVE and Pendle; 3) Withdraw, mainly from yield-farming apps; 4) Borrow, based on lending protocols like Compound; 5) Stake and 6) Claim, both predominantly from yield-farming apps like Convex Finance.
 Swap involves a user exchanging one token ($Token_{in}$) for another ($Token_{out}$) from a liquidity pool.
@@
 Borrow refers to a DeFi operation where a borrower receives a token ($Token_{borrow}$) and incurs a debt evidenced by another token ($Token_{debt}$), issued through minting.
-Stake happens in yield-farming protocols like Convex Finance (Etherscan, 2024a), where users deposit tokens and can earn rewards without receiving a minted proof token.
+Stake happens in yield-farming protocols like Convex Finance [35], where users deposit tokens and can earn rewards without receiving a minted proof token.
 Claim enables users to retrieve staked tokens ($Token_{claim}$) and bonuses without burning any tokens, unlike the Withdraw operation which requires burning tokens to retrieve assets.
@@
 To recover DeFi operations from the TG, we design and employ a search algorithm based on directed graphs for each operation.
-Since the recovery algorithm of DeFi operations is not the central focus of this work, we use the recovery of the Swap operation as an illustrative example here.
-Figure 7 illustrates that, to recover a Swap operation, we use the depth-first search algorithm to identify cycles that both start and end at user-controlled accounts.
+Fig.
+4 illustrates that, to recover a Swap operation, we use the depth-first search algorithm to identify cycles that both start and end at user-controlled accounts.
 These cycles must satisfy three constraints: (i) the transfer action of each edge must be transferring token; (ii) the time index of each edge must be monotonically increasing; and (iii) the token involved in the first transfer action must differ from the token in the last transfer action.
 Therefore, from the well-constructed TG, we can recover the Swap operation expressed as $UC\rightarrow T_{1}\rightarrow CA_{1}\rightarrow T_{4}\rightarrow CA_{2}\rightarrow T_{5}\rightarrow CA_{3}\rightarrow T_{7}\rightarrow UC$, and we label $CA_{1}$, $CA_{2}$, and $CA_{3}$ as liquidity pools.
-Table 1. Price change information-directed attack patterns, systematically mined from all types of DeFi applications (see §2.1).
-Type Pattern Details
-Buy & Sell  (BlockSec, 2024c) I 1) Swap $Token_{x}$ to $Token_{y}$ through $Pool_{buy}$ 2) The price of $Token_{y}$ in $Pool_{sell}$ increases / The price of $Token_{z}$ in $Pool_{sell}$ decreases 3) Swap $Token_{y}$ to $Token_{z}$ through $Pool_{sell}$
-II 1) The price of $Token_{x}$ in $Pool_{buy}$ increases / The price of $Token_{y}$ in $Pool_{buy}$ decreases/       The price of $Token_{y}$ in $Pool_{sell}$ increases / The price of $Token_{z}$ in $Pool_{sell}$ decreases 2) Swap $Token_{x}$ to $Token_{y}$ through $Pool_{buy}$ 3) Swap $Token_{y}$ to $Token_{z}$ through $Pool_{sell}$
-Deposit & Borrow  (Immunefi, 2024) III 1) Deposit $Token_{x}$ into $Contract_{deposit}$ and get $Token_{y}$ as credential 2) The price of $Token_{x}$ in $Contract_{borrow}$ increases / The price of $Token_{z}$ in $Contract_{borrow}$ decreases 3) Borrow $Token_{z}$ using $Token_{x}$ as collateral from $Contract_{borrow}$
-IV 1) The price of $Token_{x}$ in $Contract_{borrow}$ increases / The price of $Token_{z}$ in $Contract_{borrow}$ decreases 2) Deposit $Token_{x}$ into $Contract_{deposit}$ and get $Token_{y}$ as credential 3) Borrow $Token_{z}$ using $Token_{x}$ as collateral from $Contract_{borrow}$
-Stake & Claim  (Certik, 2024) V 1) Stake $Token_{x}$ into $Contract_{stake}$ 2) The price of $Token_{y}$ in $Contract_{claim}$ decreases 3) Claim $Token_{y}$ from $Contract_{claim}$
-VI 1) The price of $Token_{x}$ in $Contract_{stake}$ increases / The price of $Token_{y}$ in $Contract_{claim}$ decreases 2) Stake $Token_{x}$ into $Contract_{stake}$ 3) Claim $Token_{y}$ from $Contract_{claim}$
-Deposit & Withdraw  (SlowMist, 2024a) VII 1) Deposit $Token_{x}$ into $Contract_{deposit}$ and get $Token_{y}$ as credential 2) The price of $Token_{y}$ in $Contract_{withdrawal}$ increases / The price of $Token_{z}$ in $Contract_{withdrawal}$ decreases 3) Withdraw $Token_{z}$ from $Contract_{withdrawal}$ by burning $Token_{y}$
-VIII 1) The price of $Token_{x}$ in $Contract_{deposit}$ increases / The price of $Token_{y}$ in $Contract_{deposit}$ decreases /       The price of $Token_{y}$ in $Contract_{withdrawal}$ increases / The price of $Token_{z}$ in $Contract_{withdrawal}$ decreases 2) Deposit $Token_{x}$ and get $Token_{y}$ as credential 3) Withdraw $Token_{z}$ from $Contract_{withdrawal}$ by burning $Token_{y}$
-6. Price Manipulation Detection
-The price change information and high-level DeFi operations recovered from §4 and §5, respectively, are finally checked against the detection rules (c.f.
-Table 1).
-We analyzed all four specific types of DeFi protocols (c.f.
-§2.1) and their associated attack instances.
-Specifically, we examined ElephantMoney (BlockSec, 2024c) for DEX, Cream Finance (Immunefi, 2024) for lending protocols, ATK (Certik, 2024) for staking-based yield farming protocols, and Harvest (SlowMist, 2024a) for deposit-based yield farming protocols, which collectively incurred a loss of \$163.3M.
-Based on this in-depth analysis, we identify four attack types targeting different DeFi protocols and their eight generalized attack patterns, as depicted in Table 1.
+VI Price Manipulation Detection
+The price change information and high-level DeFi operations recovered from §IV and §V, respectively, are finally checked against the detection rules (c.f.
+Table I).
+We analyze four prevalent types of DeFi protocols and their associated attack instances.
+Specifically, we examine ElephantMoney [31] for DEX, Cream Finance [32] for lending protocols, ATK [33] for staking-based yield farming protocols, and Harvest [34] for deposit-based yield farming protocols, which collectively incurred a loss of \$163.3M.
+Based on this in-depth analysis, we identify four attack types targeting different DeFi protocols and their eight generalized attack patterns, as depicted in Table I.
 These attack types are Buy & Sell, Deposit & Borrow, Stake & Claim, and Deposit & Withdraw, with each type corresponding to two specific attack patterns.
-We will discuss these attack types and patterns in more detail.
-Buy & Sell.
-In this type of attack strategy, the attacker primarily profits by first buying $Token_{y}$ with $Token_{x}$ through a swap in $Pool_{buy}$ and then selling $Token_{y}$ for $Token_{z}$ through a swap in $Pool_{sell}$.
-$Token_{x}$ and $Token_{z}$ can be the same or different tokens.
-In the attack against ElephantMoney, the attacker first conducted a swap in $Pool_{buy}$ to exchange WBNB for ELEPHANT and then invoked the mint function, which triggered a swap in $Pool_{sell}$ to exchange ELEPHANT for WBNB, resulting in a price increase of ELEPHANT in $Pool_{sell}$.
-Ultimately, the attacker utilized a reverse swap in $Pool_{sell}$ to obtain WBNB by selling ELEPHANT at the manipulated price.
-We design Pattern I based on this attack and subsequently generalize it to Pattern II.
-The major difference between these two patterns is that the token price is manipulated before the first swap in Pattern II, allowing the price of tokens in either $Pool_{buy}$ or $Pool_{sell}$ to be manipulated.
-Deposit & Borrow.
-In this type, the attacker inflates the price of the deposited tokens or deflates the price of the borrowed assets as calculated by $Contract_{borrow}$, bypassing the protective mechanism of over-collateralization, thereby borrowing more assets than the actual value of the collateral.
-In the Cream Finance incident, the attacker first deposited yUSD as collateral and obtained an equivalent amount of crYUSD as proof of deposit, then inflated the price of yUSD calculated by $Contract_{borrow}$ by transferring a large quantity of yCrv to a specific contract account.
-Finally, using yUSD as collateral, the attacker borrowed a large amount of assets, which far exceeded the actual value of the deposited yUSD, from $Contract_{borrow}$.
-We design Pattern III based on this attack and then generalize it to Pattern IV.
-Pattern IV differs from Pattern III in that the attacker can preemptively increase the price of tokens designated for deposit or decrease the price of assets intended for borrowing as calculated by $Contract_{borrow}$ before the deposit operation.
-In particular, the motivating example in §2 conforms to Pattern IV.
-Stake & Claim.
-This attack type primarily targets yield-farming protocols that offer staking services.
-Typically, an attacker first stakes $Token_{x}$ into the application in one transaction.
-The share ratio of the user is calculated based on the value and quantity of the staked asset in real-time and is stored in the state variables.
-Then, the attacker decreases the calculated price of $Token_{y}$ in $Contract_{claim}$ and subsequently claims $Token_{y}$ from the contract.
-$Token_{x}$ and $Token_{y}$ can be the same or different tokens.
-We derive Pattern V based on the analysis of the attack against ATK.
-Specifically, in the first transaction, the attacker initially staked ATK into $Contract_{stake}$.
-Since the staking service required that the ATK be held for 24 hours by the contract account before claiming, the attacker waited for a period and executed the second transaction, exploiting a flash loan to deflate the price of ATK in $Contract_{claim}$, subsequently claiming back an amount of ATK significantly higher than the appropriate quantity.
-Considering that the attacker can inflate the price of tokens intended for staking beforehand to get an incorrectly calculated share ratio, we further derive Pattern VI from Pattern V.
-Deposit & Withdraw.
-In this attack type, the attacker exploits vulnerabilities in the token pricing mechanism within the deposit or withdrawal contract to conduct price manipulation attacks.
-We design and generalize Pattern VII based on the Harvest attack.
-In this hack, the attacker first deposited USDC ($Token_{x}$) in $Contract_{deposit}$ and received fUSDC ($Token_{y}$) as proof.
-Then, by exchanging USDC for USDT, the price of USDC calculated by $Contract_{withdrawal}$ decreased, and the attacker withdrew an excessive amount of USDC ($Token_{z}$) from $Contract_{withdrawal}$ by burning fUSDC.
-In this case, $Token_{x}$ and $Token_{z}$ are the same; however, some protocols, such as (Bscscan, 2024), allow different tokens for deposit and withdrawal.
-Besides deflating the price of tokens to be withdrawn, the attacker can also inflate the price of tokens used for calculating the withdrawal amount, i.e., $Token_{y}$.
-If the attacker manipulates the token price before depositing, the price of tokens involved in the deposit can also be affected.
-Based on this assumption, we generalize Pattern VII to create Pattern VIII.
-7. Implementation and Evaluation
+More details about these attack types and patterns are available in our supplementary material [28].
 We implement DeFiScope with about 3,900 lines of Python code.
-DeFiScope currently supports two blockchains, i.e., Ethereum (Ethereum, 2024) and BSC (Chain, 2024), which account for over 60% of total value locked (TVL) among all the blockchains (DeFiLlama, 2024).
-To obtain raw data from specific transactions, we utilize blockchain node APIs facilitated by QuickNode (QuickNode, 2024), an external RPC service provider.
-Our evaluation aims to answer three research questions (RQs):
+DeFiScope currently supports two blockchains, i.e., Ethereum [36] and BSC [37], which account for over 60% of total value locked (TVL) among all the blockchains [38].
+To obtain raw data from specific transactions, we utilize blockchain node APIs facilitated by QuickNode [39], an external RPC service provider.
+VII Evaluation
+We aims to evaluate the following research questions (RQs):
   • RQ1: How effectively does DeFiScope detect price manipulation attacks compared with the existing state-of-the-arts?
@@
 Datasets.
-To address these three RQs, we collect three datasets from real DeFi transactions on Ethereum and BSC to evaluate DeFiScope, as shown in Table 2.
-Specifically, we use the first dataset, D1, which comprises 95 transactions of real-world price manipulation attacks from 90 DeFi applications, to evaluate both RQ1 and RQ2.
-To demonstrate DeFiScope’s practicality for RQ3, we use the second dataset, D2, consisting of 968 suspicious transactions collected by our industry partner.
-Furthermore, to measure DeFiScope’s time overhead when deployed in a realistic setting with both suspicious and benign transactions, we mix the 968 suspicious transactions from D2 with 96,800 benign transactions in the third dataset, D3.
+To address these RQs, we collect three datasets from real DeFi transactions on Ethereum and BSC to evaluate DeFiScope, as shown in Table II.
+Specifically, we use the first dataset D1, which comprises 95 transactions of real-world price manipulation attacks from 90 DeFi applications, to evaluate both RQ1 and RQ2.
+To demonstrate DeFiScope’s practicality for RQ3, we use the second dataset D2, consisting of 968 suspicious transactions collected by our industry partner.
+Furthermore, to measure DeFiScope’s time overhead when deployed in a realistic setting with both suspicious and benign transactions, we mix 968 suspicious transactions from D2 with 96,800 benign transactions in the third dataset D3.
 For D1, we scraped data from multiple sources.
-Initially, we included all 54 price manipulation transactions identified by DeFort (Xie et al., 2024).
-We also collected all 55 price manipulation attacks documented by the renowned DeFi security GitHub repository, DeFiHackLabs (SunWeb3Sec, 2024), from October 26, 2020, to October 11, 2023.
+Initially, we included all 54 price manipulation transactions identified by DeFort [5].
+We also collected all 55 price manipulation attacks documented by the renowned DeFi security GitHub repository, DeFiHackLabs [40], from October 26, 2020, to October 11, 2023.
 To further expand our dataset, we acquired 31 transactions confirmed as price manipulation attacks from our industry partner.
 After removing duplicates among the three sources of data, we ultimately obtained 95 price manipulation attacks for D1, which collectively caused \$381.16M in losses on Ethereum and BSC.
-We also measure the distribution and monetary losses of the 95 attacks in D1 across our eight attack patterns described in §6.
-We observe that Pattern I has the highest number of cases, while Pattern III results in the greatest losses.
-This is largely due to one of the most notorious incidents, CreamFinance (Immunefi, 2024), which falls under this pattern and led to a loss of \$130M.
-For Patterns V and VIII, there is only one case each: ATK (Certik, 2024) and LUSD (Bscscan, 2024), respectively.
-Table 2. The benchmark datasets used for evaluation.
+We also measure the distribution and monetary losses of the 95 attacks in D1 across our eight attack patterns described in §VI, as shown in Table III.
+D2 comprises 968 suspicious transactions provided by our industry partner, a Web3 security company, which monitors blockchain transactions in real-time and automatically flags transactions that yield significant profits for initiators.
+All these transactions are sufficiently complex and potentially involve price manipulation.
+While manually confirming these transactions is labor-intensive and prone to errors, 155 of them have been confirmed to belong to other vulnerabilities, such as Reentrancy (OpenLeverage [41]), Unverified User Input (YIELD [42]), and Access Control Bugs (SafeMoon [43]), by developers and industry partners.
+In this way, we can stress-test DeFiScope’s detection capabilities when mixed with other types of attack transactions.
+Furthermore, to assess DeFiScope’s false alarms on benign transactions and to measure its time overhead in a realistic setting, we need to collect a large number of benign transactions, as the most majority of real-world transactions are still benign.
+In the absence of empirical data on the ratio of suspicious to benign transactions, we adopt a conservative yet reasonable ratio of 1:100, assuming one suspicious transaction for every 100 benign transactions.
+Thus, with 968 suspicious transactions in D2, we require 96,800 benign transactions for D3.
+To this end, we randomly sampled these 96,800 transactions from DeFort [5]’s dataset of 428,523 benign transactions, which includes 384,143 benign transactions on Ethereum and BSC.
+At this sample size, we achieve a 99.999% confidence level with a margin of error of 0.625%.
+TABLE II: The benchmark datasets used for evaluation.
 Dataset RQs
@@
 D3: 96,800 benign transactions RQ3
-Table 3. Summary of different attack patterns in D1.
+TABLE III: Summary of different attack patterns in D1.
 Pattern I II III IV V VI VII VIII
@@
 Loss(\$) 55.3M 70M 141.2M 43M 61K 83K 71.4M 9K
-D2 comprises 968 suspicious transactions provided by our industry partner, a Web3 security company, which monitors blockchain transactions in real-time and automatically flags transactions that yield significant profits for initiators.
-All these transactions are sufficiently complex and potentially involve price manipulation.
-While manually confirming these transactions is labor-intensive and prone to errors, 155 of them have been confirmed to belong to other vulnerabilities, such as Reentrancy (OpenLeverage (Franklin, 2024)), Unverified User Input (YIELD (BlockSec, 2024e)), and Access Control Bugs (SafeMoon (zokyo, 2024)), by developers and industry partners.
-In this way, we can stress-test DeFiScope’s detection capabilities when mixed with other types of attack transactions.
-Furthermore, to assess DeFiScope’s false alarms on benign transactions and to measure its time overhead in a realistic setting, we need to collect a large number of benign transactions, as the most majority of real-world transactions are still benign.
-In the absence of empirical data on the ratio of suspicious to benign transactions, we adopt a conservative yet reasonable ratio of 1:100, assuming one suspicious transaction for every 100 benign transactions.
-Thus, with 968 suspicious transactions in D2, we require 96,800 benign transactions for D3.
-To this end, we randomly sampled these 96,800 transactions from DeFort (Xie et al., 2024)’s dataset of 428,523 benign transactions, which includes 384,143 benign transactions on Ethereum and BSC.
-At this sample size, we achieve a 99.999% confidence level with a margin of error of 0.625%.
+Minimized Data Leakage.
+DeFiScope utilizes LLMs to assess token price changes solely by using price calculation functions and token balance changes, without any knowledge of attack transactions.
+Even if related attacks appeared in pre-training, evaluating price change tendencies in this manner falls outside its training data, resulting in no or minimized data leakage.
 Experimental Setup.
 All experiments were conducted on a desktop computer running Ubuntu 20.04, powered by an Intel® Xeon® W-2235 CPU (3.80 GHz, 6 cores, and 12 threads) and equipped with 16 GB of memory.
-For LLMs, we use OpenAI’s GPT family models due to limited data required for fine-tuning, as explained in §4.1.
-DeFiScope by default uses GPT-3.5-Turbo (GPT-3.5-turbo-1106) for its well-recognized cost-performance balance, but we also conduct an ablation study in §7.2 to test the more advanced GPT-4o (GPT-4o-2024-08-06) for fine-tuning.
-For the LLM configuration, DeFiScope employs nucleus sampling (Holtzman et al., 2019) with a top-p value of 1 and sets the temperature to 0, ensuring highly deterministic responses for each prompt, although each was run only once.
-Table 4.
-Detection results for 95 ground-truth DeFi price manipulation attacks.
-DS: DeFiScope, DT: DeFiTainter, DF: DeFort, and DR: DeFiRanger.
-DeFiScope detected 76 attacks, DeFiTainter detected 34 attacks, DeFort detected 50 attacks, and DeFiRanger detected 49 attacks.
-Note that DeFiScope failed in 8 cases solely due to ① missing source code and ② compilation errors, which were not caused by DeFiScope’s methodology.
+For LLMs, we by default use OpenAI’s GPT family models but also explore open-source SFT model tuning in §VIII, as explained in §IV-A.
+DeFiScope by default uses GPT-3.5-Turbo (GPT-3.5-turbo-1106) for its well-recognized cost-performance balance, but we also conduct an ablation study in §VII-B to test the more advanced GPT-4o (GPT-4o-2024-08-06) for fine-tuning.
+For the LLM configuration, DeFiScope employs nucleus sampling [44] with a top-p value of 1 and sets the temperature to 0, ensuring highly deterministic responses for each prompt, although each was run only once.
+TABLE IV: Detection results for 95 ground-truth DeFi price manipulation attacks.
 Protocol Chain Date Loss DS DT DF DR
@@
 Zunami ETH 13-Aug-23 2M
-7.1. RQ1: Detection Effectiveness
-To answer RQ1, we evaluate DeFiScope and compare it with three SOTA tools—DeFiTainter (Kong et al., 2023), DeFort (Xie et al., 2024), and DeFiRanger (Wu et al., 2023), using the dataset D1.
-Because DeFiRanger is not open-source, we use the results reported in their paper for the attacks they evaluated; for other attacks, we re-implemented their approach based on the description (Wu et al., 2023) and conducted our evaluation.
-Although DeFort is also not open-source, we obtained a copy of its source code from the authors to conduct our experiment.
-Table 4 presents the detection results.
-The first four columns list the name of the victim protocol, the destination chain, the hack date, and the resulting loss, respectively.
+VII-A RQ1: Detection Effectiveness
+To answer RQ1, we evaluate DeFiScope and compare it with three SOTA tools, DeFiTainter [7], DeFort [5], and DeFiRanger [4], specialized in detecting price manipulations using the dataset D1.
+Because DeFiRanger is not open-source, we use the results reported in their paper for the attacks they evaluated; for other attacks, we re-implemented their approach based on the description [4] and conducted our evaluation.
+Although DeFort is also not open-source, we obtained a copy of code from its authors to run experiments.
+FlashSyn [45] and SMARTCAT [46] were excluded as baselines because FlashSyn targets broader attacks beyond price manipulations and lacks guidelines for detecting new attacks, while SMARTCAT was unpublished and not open-source at the time of evaluation.
+Table IV presents the detection results.
+The first four columns list the name of the victim protocol, chain deployed, hack date, and resulting loss, respectively.
+DS, DT, DF and DR denote DeFiScope, DeFiTainter, DeFort and DeFiRanger, respectively.
 Note that one protocol may face multiple attacks, so we add numerical suffixes to protocol names to differentiate them.
@@
 Overall, DeFiScope detected 76 attacks, followed by DeFort with 50 and DeFiRanger with 49 attacks, respectively, while DeFiTainter detected only 34 attacks.
-Figure 8 details the performance of each tool on the evaluated DeFi protocols across four application categories using different price models.
+In total, DeFiScope failed to detect 19 cases.
+We manually investigated their root cause and found: (a) 11 cases should be detectable but were missed by DeFiScope due to the limitation of LLM reasoning, non-ERC20 token and the restriction to only single-transaction analysis, and (b) 8 cases cannot be detected because they do not satisfy the source code requirement of DeFiScope due to either ① missing source code or ② compilation errors, as marked in Table IV.
+Moreover, the statistical significance of DeFiScope’s performance improvements over each baseline is confirmed by the McNemar Test [47], with all the resulting $p$-values being below 0.05 (see details in  [28]).
+Fig.
+5 details the performance of each tool on the evaluated DeFi protocols across four application categories using different price models.
 Compared to other tools, DeFiScope performs the best in every application category.
-Particularly, it achieves the highest recall, 90.7%, in Token-related protocols.
+Particularly, it achieves the highest recall, 90.7%, in the Token category.
 However, DeFiScope yields a low recall rate in Lending-related protocols, though still higher than all existing tools.
@@
 Additionally, one attack (TIFIToken) involves exploiting a closed-source custom price model, and the InverseFinance_1 attack, a false negative, will be detailed below.
-Yet, DeFiScope missed detecting 11 price manipulation attacks.
-After analyzing each attack, we discovered that 8 out of 11 are cross-transaction attacks (Certik, 2024; Ancilia, 2024b; Phalcon, 2024; Inspex, 2024; Knownsec, 2024; Ancilia, 2024c; BlockSec, 2024d; Finance, 2024), where detection was unsuccessful because DeFiScope is based on analyzing individual transactions.
+After analyzing 11 detectable but missed attacks, we discovered that 8 out of 11 are cross-transaction attacks [33, 48, 49, 50, 51, 52, 53, 54], where detection was unsuccessful because DeFiScope is based on analyzing individual transactions.
 For Zoompro, detection failure occurred because the token involved in the transaction did not adhere to the ERC20 token standard, resulting in the transfer event not being identified.
@@
 As both cases require precise quantitative calculations, DeFiScope is limited by the current LLMs’ constrained capacity for scientific computation.
-A potential solution could be to integrate with Program-Aided Language models (PAL) (Gao et al., 2023), guiding the LLM to generate scripts for necessary calculations and executing them to obtain the result.
-Among the remaining 8 undetectable transactions, 3 were not analyzed successfully due to the unavailability of the code of price calculation functions since our method relies on code-level analysis; five cases involved compilation errors during the extraction of price calculation functions using Slither (crytic, 2024), an off-the-shelf static analyzer for Solidity and Vyper, thereby prematurely terminating the detection process.
-These limitations are not inherent to DeFiScope’s methodology and can be mitigated by automated or semi-automated techniques, e.g., code decompilation (Grech et al., 2019, 2022; Suiche, 2017) and manual intervention.
-Figure 8. The categorized detection results on DeFi protocols.
-7.2. RQ2: Ablation Study
+A potential solution could be to integrate with Program-Aided Language models (PAL) [55], guiding the LLM to generate scripts for necessary calculations and executing them to obtain the result.
+Among the 8 undetectable transactions, 3 were not analyzed successfully due to the unavailability of the code of price calculation functions since our method relies on code-level analysis; five cases involved compilation errors during the extraction of price calculation functions using Slither [56], an off-the-shelf static analyzer for Solidity and Vyper, thereby prematurely terminating the detection process.
+These limitations are not inherent to DeFiScope’s methodology and can be mitigated by automated or semi-automated techniques, e.g., code decompilation [57, 58, 59] and manual intervention.
+Figure 5: The categorized detection results on DeFi protocols.
+VII-B RQ2: Ablation Study
 In this RQ, we investigate how fine-tuning can enhance DeFiScope’s detection accuracy on the same D1 dataset, as well as the impact and cost of fine-tuning different GPT models.
-To this end, we test four settings shown in Figure 9: (a) the original GPT-3.5-Turbo without fine-tuning, (b) GPT-3.5-Turbo with fine-tuning, which was used in RQ1, (c) the original GPT-4o without fine-tuning, and (d) GPT-4o with fine-tuning.
-Figure 9 shows that fine-tuning significantly enhances DeFiScope’s ability to detect attacks, with the fine-tuned versions of GPT-3.5-Turbo and GPT-4o detecting 18 (31%) and 12 (19%) more attacks, respectively.
+To this end, we test four settings shown in Fig.
+6: (a) the original GPT-3.5-Turbo without fine-tuning, (b) GPT-3.5-Turbo with fine-tuning, used in RQ1, (c) the original GPT-4o without fine-tuning, and (d) GPT-4o with fine-tuning.
+Fig.
+6 shows that fine-tuning significantly enhances DeFiScope’s ability to detect attacks, with the fine-tuned versions of GPT-3.5-Turbo and GPT-4o detecting 18 (31%) and 12 (19%) more attacks, respectively.
 It also indicates that fine-tuning provides a more noticeable improvement for less powerful models, such as GPT-3.5-Turbo, compared to stronger models like GPT-4o.
-With fine-tuning, the detection success rate for attacks targeting the CPMM increases to 100% with GPT-3.5-Turbo and 95.6% with GPT-4o [^4] GPT-4o does not exhibit an advantage with CPMM but shows a clear advantage with custom price models, especially when the original GPT-4o compared to GPT-3.5-Turbo..
-This very high rate can be attributed to the use of CPMM during the fine-tuning training phase, enabling the model to effectively handle this specific pricing model.
+With fine-tuning, the detection success rate for attacks targeting CPMM increases to 100% with GPT-3.5-Turbo and 95.6% with GPT-4o [^2] GPT-4o does not exhibit an advantage with CPMM but shows a clear advantage with custom price models, especially for the original GPT-4o..
+This very high rate can be attributed to the use of CPMM during the fine-tuning training phase, enabling the model to effectively handle it.
 (a) GPT-3.5-Turbo w/o fine-tuning
@@
 (d) GPT-4o w/ fine-tuning
-Figure 9.
-Effectiveness of fine-tuning for 78 attacks that go through DeFiScope’s LLM inference.
-The cases not plotted in the figures include eight due to missing source code and compilation errors, and nine due to cross-transaction issues and non-adherence to the ERC20 token standard, as mentioned in §7.1.
+Figure 6: Effectiveness of fine-tuning for 78 attacks that go through DeFiScope’s LLM inference.
+Cases not plotted in the figures include eight due to missing source code and compilation errors, and nine due to cross-transaction issues and non-adherence to the ERC20 token standard, as mentioned in §VII-A.
 Although DeFiScope’s fine-tuned models were trained using only CPMM data, we find that they also exhibit strong transfer learning capabilities for attacks targeting custom price models.
@@
 In contrast, although the original model produces the final evaluation scores, it does not strictly follow the instructions of the CoT prompts.
-Table 5. Comparison between GPT-3.5-Turbo and GPT-4o.
-Model $TP$ $Recall$ Ave. Cost(\$)/ per inference Fine-tuning Cost(\$)
+LLM Costs.
+Table V highlights that while the original GPT-4o generally performs better than GPT-3.5-Turbo, this performance gap can be largely minimized through fine-tuning.
+However, regarding costs, GPT-3.5-Turbo has a clear advantage, which is why DeFiScope uses GPT-3.5-Turbo for fine-tuning by default instead of GPT-4o.
+Specifically, the fine-tuning cost of GPT-3.5-Turbo is 68% lower than that of GPT-4o, and the per-request inference cost of GPT-3.5-Turbo is merely \$0.0107 [60], which is also 18% cheaper than that of GPT-4o.
+Therefore, for a tradeoff between cost and performance, we recommend using GPT-3.5-Turbo.
+TABLE V: Comparison between GPT-3.5-Turbo and GPT-4o.
+Model TPTP RecallRecall Ave. Cost(\$)/ per inference Fine-tuning Cost(\$)
 GPT-3.5 w/o fine-tuning 58 0.61 0.0023 -
@@
 GPT-4o w/ fine-tuning 75 0.79 0.0131 25
-LLM Costs.
-Table 5 highlights that while the original GPT-4o generally performs better than GPT-3.5-Turbo, this performance gap can be largely minimized through fine-tuning.
-However, regarding costs, GPT-3.5-Turbo has a clear advantage, which is why DeFiScope uses GPT-3.5-Turbo for fine-tuning by default instead of GPT-4o.
-Specifically, the fine-tuning cost of GPT-3.5-Turbo is 68% lower than that of GPT-4o, and the per-request inference cost of GPT-3.5-Turbo is merely \$0.0107 (OpenAi, 2024b), which is also 18% cheaper than that of GPT-4o.
-Therefore, for a tradeoff between cost and performance, we recommend using GPT-3.5-Turbo for DeFiScope.
-7.3. RQ3: Real-World Practicality
+VII-C RQ3: Real-World Practicality
 To be practical in a real-world setting, DeFiScope not only needs to maintain high detection rates for true attacks but also minimize false alarms for benign transactions.
-In this RQ, we evaluate this aspect of DeFiScope and the associated time overhead using the datasets D2 and D3, which were introduced in the prologue of §7.
-Specifically, D2 comprises 968 suspicious transactions with various attacks (i.e., not limited to price manipulations), and D3 includes 96,800 benign transactions collected from DeFort (Xie et al., 2024)’s dataset.
+In this RQ, we evaluate this aspect of DeFiScope and the associated time overhead using the datasets D2 and D3, which were introduced in the prologue of §VII.
+Specifically, D2 comprises 968 suspicious transactions with various attacks (i.e., not limited to price manipulations), and D3 includes 96,800 benign transactions collected from DeFort [5]’s dataset.
 For D2, DeFiScope flagged 153 price manipulation attacks out of 968 suspicious transactions.
-To robustly confirm these potential attacks, we cross-referenced them with attack reports or alerts published by security companies through their official channels (Ancilia, 2024a; Beosin, 2024; BlockSec, 2024b; SlowMist, 2024b), thereby verifying the root causes of the attacks.
+To robustly confirm these potential attacks, we cross-referenced them with attack reports or alerts published by security companies through their official channels [61, 62, 63, 64], thereby verifying the root causes of the attacks.
 Using this method, we confirm that 66 of them are previously reported price manipulation attacks.
-For the remaining cases, we conducted comprehensive and in-depth analyses by combining manual review with ancillary evidence, such as identifying whether the EOA initiating the transaction was marked as a hacker by blockchain explorers (BscScan, 2024; EtherScan, 2024; OKLink, 2024).
-Finally, we discovered 81 previously unknown historical incidents, which were neither reported by security companies nor tagged as malicious transactions by blockchain explorers, and identified a total of six false positive cases (as detailed in Table 6 in Appendix C).
-Of these, five are benign transactions initiated by the same EOA and exhibiting identical invocation flows, while the last one is indeed an attack due to a logic issue rather than a price manipulation attack, resulting in a precision of 96% (147/153).
+For the remaining cases, we conducted comprehensive and in-depth analysis by combining manual review with ancillary evidence, such as identifying whether the EOA initiating the transaction was marked as a hacker by blockchain explorers [65, 66, 67].
+For newly discovered incidents, we manually verified: (i) whether abnormal token price or exchange rate changes occured, and (ii) whether the user gained profit from such changes.
+Cases satisfying both criteria were labeled as price manipulation attacks.
+Finally, we discovered 81 previously unknown historical incidents, which were neither reported by security companies nor tagged as malicious transactions by blockchain explorers, and identified a total of six false positive cases (the full list can be found in [28]).
+Of these, five are benign transactions initiated by the same EOA and exhibiting identical invocation flows, while the last one is indeed an attack due to a logic issue rather than a price manipulation attack, resulting in a precision of 96% (147/153) on suspicious transactions.
 In comparison, DeFort identified only 58 attacks, of which 9 were false positives labeled as other vulnerabilities by security researchers.
@@
 For each transaction, we set the maximum scan time to 300 seconds.
-The results reveal that DeFiScope achieves zero false alarms on benign transactions in D3, with an average of 2.5 seconds per transaction across all transactions.
+The results reveal that DeFiScope achieves zero false alarms on benign transactions in D3, and 6 false positives on D2 (as mentioned above), with an average of 2.5 seconds per transaction across all transactions.
+More than half the time is spent on price change inference that uses up an average of 1.40 seconds, while static analysis takes slightly less, around 0.97 seconds on average.
+The other steps within DeFiScope incur negligible cost.
+A detailed time cost breakdown is available in  [28].
 This highlights DeFiScope’s potential for large-scale, daily on-chain monitoring scenarios.
-8. Discussion
-While the precision of DeFiScope is quite satisfactory, its recall rate primarily affected by the following two threats. Future work can further optimize the following two aspects:
-Cross-transaction price manipulation attacks.
-DeFiScope currently detects only single-transaction attacks.
-However, some price manipulation attacks are designed across multiple transactions to circumvent time restrictions coded in DeFi protocols, as studied in (Chen et al., 2024b).
-Detecting cross-transaction attacks is often challenging.
-For example, attacks like INUKO (Ancilia, 2024b) span 48 hours (crossing around 57,000 blocks), making it extremely difficult to precisely identify all related transactions.
-Fortunately, our study observed that cross-transaction attacks are less common than single-transaction ones.
-Closed-source price calculation functions.
-The availability of code for price models could affect the detection accuracy of DeFiScope.
+VIII Discussion
+Cross-LLM Generalizability.
+To show DeFiScope’s methodology is not limited to OpenAI’s fine-tuning paradigm [22], we performed SFT fine-tuning [23] on a open-source Phi-3 [68] model using LoRA [24], on a sever with 4 Nvidia H800 GPUs.
+We used all 1,000 price change samples mentioned in §IV, split 8:1:1 into training, evaluation, and test sets.
+As shown in Table VI, LoRA-based fine-tuning enabled Phi-3 to outperform GPT-3.5-Turbo (without fine-tuning) and achieve performance comparable to GPT-4o (without fine-tuning).
+Full hyperparameter settings and results are available in our repository [69].
+TABLE VI: Comparison between GPT tuning and Phi-3’s SFT.
+Model RecallRecall Model RecallRecall
+GPT-3.5 w/o fine-tuning 0.61 GPT-3.5 w/ fine-tuning 0.80
+GPT-4o w/o fine-tuning 0.66 GPT-4o w/ fine-tuning 0.79
+Phi-3 w/o fine-tuning 0.52 Phi-3 w/ fine-tuning [70] 0.66
+Cross-Blockchain Generalizability.
+Although our evaluation primarily focused on Ethereum and BSC, which together account for the majority of observed price manipulation attacks, DeFiScope is generalizable to EVM-compatible blockchains, such as Polygon [71] and Arbitrum [72].
+DeFiScope successfully detected seven price manipulation attacks on these chains.
+Full details of these cases are available in [28].
+Closed-source contracts and cross transaction attacks.
+The availability of price model code can affect DeFiScope’s detection accuracy.
 According to our study, most DeFi applications are open-source to gain user trust.
 To mitigate issues with closed-source price models, we design the Type-II prompt to cover price models in those closed-source liquidity pools.
-9. Related Work
+Currently, DeFiScope detects only single-transaction attacks, while some attacks span multiple transactions to bypass protocol time restrictions [73].
+Detecting such cases is challenging.
+For instance, attacks like [48] span over 48 hours (about 57,000 blocks), making it difficult to trace all related transactions.
+Fortunately, our study observed that such attacks are far less common than single-transaction ones.
+IX Related Work
 On-chain Security Analysis.
-Prior efforts (Zhou et al., 2021b; Qin et al., 2021, 2022; Daian et al., 2020; Zhou et al., 2023) have elucidated threats in blockchain and attacks against the DeFi ecosystem.
-Some transaction-based systems have been proposed to mine vulnerable transaction sequences (Zhang et al., 2020; So et al., 2021), explore arbitrage opportunities (Zhou et al., 2021a), detect malicious phishing (He et al., 2023), and simulate attacks to prevent intrusions (Qin et al., 2023a).
-Tools like FlashSyn (Chen et al., 2024a), DeFiRanger (Wu et al., 2023), DeFiTainter (Kong et al., 2023), and DeFort (Xie et al., 2024) are capable of detecting attacks associated with price manipulation.
-FlashSyn uses numerical approximation techniques to synthesize malicious contracts that target DeFi apps through price manipulation attacks.
-However, there is a gap between the synthesized contracts and those that launch attacks on the live blockchain.DeFiRanger generates cash flow trees (CFT) from a transaction, translates the semantics of CFT into high-level DeFi actions, and uses specific patterns to detect price manipulation attacks.
-Yet, DeFiRanger is significantly limited by its reliance on predefined templates.
-DeFiTainter constructs call graphs based on contract states and transactions, then performs inter-contract taint analysis to examine all execution paths and uncover price manipulation attacks.
-Nonetheless, DeFiTainter’s approach requires manual analysis and labeling of taint sources and sinks, which is challenging.
-DeFort uses a price manipulation behavior model to guide detection and analysis but its boundary calculation mechanism lacks sensitivity to attacks characterized by minor price fluctuations.
-Program Analysis for Contract Security.
-Traditional tools such as Slither (Feist et al., 2019), Zeus (Kalra et al., 2018), and Securify (Tsankov et al., 2018) employ static analysis methods to detect vulnerabilities at the source code and bytecode levels.
-Symbolic execution has also shown significant potential in this area, with tools like Mythril (Consensys, 2024), Oyente (Luu et al., 2016), and Manticore (Mossberg et al., 2019) developed based on this technique.
-Dynamic analysis is another core methodology utilized by various tools (Jiang et al., 2018; Liu et al., 2018; Shou et al., 2023; Qin et al., 2023b).
-Some tools are specifically designed to detect particular types of vulnerabilities; for example, VetSC (Duan et al., 2022) identifies business logic bugs by vetting inconsistencies between UI and contract code, SailFish (Bose et al., 2022) focuses on state inconsistency issues, and JACKAL (Gritti et al., 2023) targets confused deputy vulnerabilities.
+Prior efforts [74, 75, 76, 77, 78] have studied blockchain threats and DeFi attacks.
+Some transaction-based systems are proposed to mine vulnerable transaction sequences [79, 80], explore arbitrage opportunities [81], detect malicious phishing [82], and simulate attacks to prevent intrusions [83].
+Tools like FlashSyn [45], DeFiRanger [4], DeFiTainter [7], and DeFort [5] are capable of detecting attacks associated with price manipulation.
+FlashSyn uses numerical approximation to synthesize malicious contracts that target DeFi apps through price manipulation attacks.
 LLMs for Smart Contract Security.
-LLMs have become formidable tools in the realm of blockchain security.
-GPTScan (Sun et al., 2024b) uses LLMs for code understanding, incorporating static analysis to detect logical vulnerabilities in contracts.
-BlockGPT (Gai et al., 2023) serves as a real-time intrusion detection system, using LLMs to rank transaction anomalies.
-LLM4Vuln (Sun et al., 2024a) refines the reasoning processes of LLMs when analyzing smart contract vulnerabilities.
-iAudit (Ma et al., 2025) combines fine-tuning with LLM-driven agents to provide intuitive smart contract auditing with explanations.
-PropertyGPT (Liu et al., 2025) uses retrieval-augmented property generation for smart contracts, leveraging LLMs’ in-context learning capabilities.
-Similarly, SMARTINV (Wang et al., 2024) employs Tier of Thought (ToT) to reason across different modalities of smart contracts, generating invariants in the process.
-To the best of our knowledge, DeFiScope is the first tool that uses LLMs specifically designed for detecting price manipulation attacks.
-10. Conclusion
+LLMs have become powerful tools in blockchain security.
+GPTScan [84] leverages LLMs with static analysis to detect logical vulnerabilities.
+BlockGPT [85] uses LLMs to rank transaction anomalies for real-time intrusion detection.
+LLM4Vuln [86] improves LLM reasoning for vulnerability analysis.
+iAudit [25] combines fine-tuning with agents for intuitive auditing.
+PropertyGPT [87] uses LLMs for retrieval-augmented property generation.
+To our knowledge, DeFiScope is the first tool that uses LLMs specifically designed for detecting price manipulation attacks.
+X Conclusion
 In this paper, we introduced DeFiScope, the first tool that utilizes LLMs specifically for detecting price manipulation attacks.
@@
 Future work includes better handling of closed-source price calculation functions.
-Appendix
-Table 6. Details of six false positives mentioned in §7.3.
-Transaction hash Chain Block Type Root Cause
-0x130c6370 BSC 38218540 Benign -
-0x4b59af93 BSC 38218538 Benign -
-0xe158a2b9 BSC 38218537 Benign -
-0x59942848 BSC 38218536 Benign -
-0x2e9ceb16 BSC 38218539 Benign -
-0x640ce34c BSC 11403670 Malicious Logic issue
-Appendix A A Case Study of Type-II Prompt
-Figure 10 illustrates an use case of Type-II prompt.
-The contract (0x2120…3379) is closed source DEX contract that allows users to trade SVT and USDT.
-While recovering the DeFi operations, this contract was identified as a two-token liquidity pool.
-So, we applied the Type-II prompt to inferring the token price trend in its transactions.
-As shown by the LLM response in Figure 10, the primary difference from the response to a Type-I prompt lies in the absence of the analysis on the price model; instead, the scoring is directly yielded.
-The is because the prompt itself assumes that the CPMM is employed in this contract.
-Regarding the CPMM, the relationship between token price and token balance is quite standardized, i.e., the direction of balance change is opposite to that of price change.
-Figure 10. An example of Type-II prompt and its response generated while detecting the attack against SVT (BlockSec, 2024a).
-Table 7.
-High-value DeFi applications (from DefiLlama (DefiLlama, 2024)) with active transactions of each category we have done a comprehensive study on.
-TVL refers to the Total Value Locked in the app.
-DEX App TVL Lending App TVL Yield farming App TVL
-Uniswap \$4.8B Compound \$2.0B Pendle \$2.8B
-Curve \$1.9B AAVE \$12.3B Convex Finance \$1.1B
-Balancer \$777.2M Morpho \$1.5B Aura \$361.9M
-Sushi \$250.2M Fraxlend \$134.4M Magpie \$192.5M
-PancakeSwap \$1.7B Venus \$1.5B StakeDAO \$78.8M
-1inch \$4.58M Strike \$9.7M Equilibria Finance \$80.0M
-ParaSwap \$6.34M Planet \$1.2M Kine Finance \$8.0M
-ShibaSwap \$18.19M Kinza Finance \$42.8M Dot Dot Finance \$2.1M
-BiSwap \$27.1M Radiant \$7.8M Solo Top \$1.8M
-MDEX \$16.0M Ambit \$6M Jetfuel Finance \$1.5M
-Appendix B High-value DeFi Applications
-The design of DeFi operations and the category of transfer actions in §5 is based on an in-depth study of the high-value DeFi applications listed in Table 7 as of August 2024.
-Appendix C Details of False Positives
-Table 6 presents the details of six false positives discovered in our experiments described in §7.3.
-The reason for these false positives is that their transactions involve two contract accounts that were created three months ago by the transaction initiator.
-Any fund transfers among these accounts and the initiator should be considered benign operations rather than price manipulation operations.
-Yet, these accounts were incorrectly marked as closed-source DEXes in the detection, leading to false inferences.
-Such false positives could be mitigated by conducting a historical analysis of account ownership relationships and clustering user accounts that are controlled by the same owners.
+[…]
+Acknowledgment We thank all reviewers for their constructive feedback.
+This research was supported by Lingnan Grant SUG-002/2526, HKUST TLIP Grant FF612, the National Natural Science Foundation of China (Project No.
+72304232), the Singapore Ministry of Education Academic Research Fund Tier 2 (T2EP20224-0003) and the Nanyang Technological University Centre for Computational Technologies in Finance (NTU-CCTF).
+Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of MOE and NTU-CCTF.